### v3

# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

Abstract
The availability of open-source molecular simulation software packages allows scientists and engineers to focus on running and analyzing simulations without having to write, parallelize, and validate their own simulation software. While molecular simulations thus become accessible to a larger audience, the “black box” nature of such software packages and wide array of options and features can make it challenging to use them correctly, particularly for beginners in the topic of simulations. LAMMPS is one such versatile molecular simulation code, designed for modeling particle-based systems across a broad range of materials science and computational chemistry applications, including atomistic, coarse-grained, mesoscale, grid-free continuum, and discrete element models. LAMMPS is capable of efficiently running simulations of varying sizes from small desktop computers to large-scale supercomputing environments. Its flexibility and extensibility make it ideal for complex and extensive simulations of atomic and molecular systems, and beyond. This article introduces a suite of tutorials designed to make learning LAMMPS more accessible to new users. The first four tutorials cover the basics of running molecular simulations in LAMMPS with systems of varying complexities. The second four tutorials address more advanced molecular simulation techniques, specifically the use of a reactive force field, grand canonical Monte Carlo, enhanced sampling, and the REACTER protocol. In addition, we introduce LAMMPS–GUI, an enhanced cross-platform graphical text editor specifically designed for use with LAMMPS and able to run LAMMPS directly on the edited input. LAMMPS–GUI is used as the primary tool in the tutorials to edit inputs, run LAMMPS, extract data, and visualize the simulated systems.
1 Introduction
Molecular simulations can be used to model a large variety of atomic and coarse-grained systems, including solids, fluids, polymers, and biomolecules, as well as complex interfaces and multi-component systems. While various molecular modeling methods exist, molecular dynamics (MD) and Monte Carlo (MC) are most commonly used. MD is the preferred method for obtaining the accurate dynamics of a system, as it relies on solving Newton’s equations of motion. For systems with many degrees of freedom or complex energy landscapes, the MC method can be a better choice than MD because it allows for efficiently exploring a configuration space without being confined by the accessible time scale. MC involves performing random changes to the system configuration that are either accepted or rejected based on energy criteria [1, 2]. Molecular simulations allow for measuring a broad variety of properties, including structural properties (e.g., bond length distribution, coordination numbers, radial distribution functions), thermodynamic properties (e.g., temperature, pressure, volume), dynamic behaviors (e.g., diffusion coefficient, viscosity), and mechanical responses (e.g., elastic constant, Poisson ratio). Some of these quantities can be directly compared with experimental data, enabling the validation of the simulated system, while others, available only through simulations, are often useful for interpreting experimental data [3].
LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulator) [4] is a highly flexible and parallel open-source molecular simulation tool. Over the years, a broad variety of particle interaction models have been implemented in LAMMPS, enabling it to model a wide range of systems, including atomic, polymeric, biological, metallic, reactive, granular, mesoscale, grid-free continuum, and coarse-grained systems [5]. LAMMPS can be used on a single CPU core, a multi-socket and multi-core server, an HPC cluster, or a high-end supercomputing system. It can efficiently handle complex and large-scale simulations, including hybrid MPI-OpenMP parallelization and MPI + GPU acceleration (for a subset of its functionality).
LAMMPS requires users to write detailed input files, a task that can be particularly challenging for new users. Although its documentation extensively describes all available features [6], navigating it can be challenging. Much of the information may be unnecessary for common use cases, and the detailed manual can often feel overwhelming. Beyond the intrinsic complexity of LAMMPS, performing accurate simulations requires several complex, system-specific decisions regarding the physics to be modeled, such as selecting the thermodynamic ensemble (e.g., microcanonical, grand canonical), determining the simulation duration, and choosing the sets of parameters describing the interactions between atoms (the so-called force field) [7, 8, 9]. While these choices are independent of the simulation software, they may occasionally be constrained by the features available in a given package. The tutorials in this article aim to flatten the learning curve and guide users in performing accurate and reliable molecular simulations with LAMMPS.
1.1 Scope
This set of tutorials consists of eight tutorials arranged in order of increasing difficulty. Although each tutorial can be read independently, information introduced in earlier tutorials is generally not repeated in detail in later ones. For this reason, we recommend that beginners follow the tutorials in order. The novelties associated with each tutorial are briefly described below.
In Tutorial 1, the structure of LAMMPS input files is illustrated through the creation of a simple atomic Lennard-Jones fluid system. Basic LAMMPS commands are used to set up interactions between atoms, perform an energy minimization, and finally run a simple MD simulation in the microcanonical (NVE) and canonical (NVT) ensembles.
In Tutorial 2, a more complex system is introduced in which atoms are connected by bonds: a small carbon nanotube. The use of both classical and reactive force fields (here, OPLS-AA [10] and AIREBO [11], respectively) is illustrated. An external deformation is applied to the CNT, and its deformation is measured. This tutorial also demonstrates the use of an external tool to visualize breaking bonds, and show the possibility to import LAMMPS-generated YAML log files into Python.
In Tutorial 3, two components—liquid water (flexible three-point model) and a polymer molecule—are merged and equilibrated. A long-range solver is used to handle the electrostatic interactions accurately, and the system is equilibrated in the isothermal-isobaric (NPT) ensemble; then, a stretching force is applied to the polymer. Through this relatively complex solvated polymer system, the tutorial demonstrates how to use type labels to make molecule files more generic and easier to manage [12].
In Tutorial 4, an electrolyte is confined between two walls, illustrating the specifics of simulating systems with fluid-solid interfaces. With the rigid four-point TIP4P/2005 [13] water model, this tutorial uses a more complex water model than Tutorial 3. A non-equilibrium MD is performed by imposing shear on the fluid through moving the walls, and the fluid velocity profile is extracted.
In Tutorial 5, the ReaxFF reactive force field, which is specifically designed to simulate chemical reactions by dynamically adjusting atomic interactions [14], is used. ReaxFF includes charge equilibration (QEq), a method that allows the partial charges of atoms to adjust according to their local environment.
In Tutorial 6, the adsorption of a fluid in silica pores is modeled. To do so, a Monte Carlo simulation in the grand canonical ensemble is implemented to demonstrate how LAMMPS can be used to simulate an open system that exchanges particles with a reservoir.
In Tutorial 7, an advanced free energy method called umbrella sampling is implemented. By calculating an energy barrier, this tutorial describes a protocol for addressing energy landscapes that are difficult to sample using classical MD or MC methods.
In Tutorial 8, a CNT embedded in nylon-6,6 polymer melt is simulated. The REACTER protocol is used to model the polymerization of nylon, and the formation of water molecules is tracked over time [15].
2 Prerequisites
2.1 Background knowledge
This set of tutorials assumes no prior knowledge of the LAMMPS software itself. To complete the tutorials, a text editor and a suitable LAMMPS executable are required. We use LAMMPS–GUI [16] here, as it offers features that make it particularly convenient for tutorials, but other console or graphical text editors, such as GNU nano, vi/vim, Emacs, Notepad, Gedit, and Visual Studio Code can also be used. LAMMPS can be executed either directly from LAMMPS–GUI (Appendix A) or from a command prompt (Appendix B), the latter of which requires some familiarity with executing commands from a terminal or command-line prompt.
In addition, prior knowledge of the theoretical basics of molecular simulations and statistical physics is highly beneficial. Users may refer to textbooks such as Understanding Molecular Simulation by Daan Frenkel and Berend Smit [1], as well as Computer Simulation of Liquids by Michael Allen and Dominic Tildesley [2]. To better understand the fundamental concepts behind the soft matter systems simulated in these tutorials, users can also refer to Basic Concepts for Simple and Complex Liquids by Jean-Louis Barrat and Jean-Pierre Hansen [17], as well as Theory of Simple Liquids: with Applications to Soft Matter by Jean-Pierre Hansen and Ian Ranald McDonald [18]. For more resources, the SklogWiki platform provides a wide range of information on statistical mechanics and molecular simulations [19].
2.2 Software/system requirements
The LAMMPS stable release version 22Jul2025 [20] and the matching LAMMPS–GUI software version 1.7.0 are required to follow the tutorials, as they include features that were first introduced in these versions. For Linux (x86_64 CPU), macOS (BigSur or later), and Windows (10 and 11) you can download a pre-compiled LAMMPS package from the LAMMPS release page on GitHub [21]. Select a package with ‘GUI’ in the file name, which includes both, LAMMPS–GUI and the LAMMPS command-line executable. These pre-compiled packages are designed to be portable, and therefore omit support for parallel execution with MPI. Instructions for installing LAMMPS–GUI and using its most relevant features for the tutorials are provided in Appendix A.
LAMMPS versions are generally backward compatible, meaning that older input files typically work the same with newer versions of LAMMPS. However, forward compatibility is not as strong, so input files written for a newer version may not always work with older versions. As a result, it is usually possible to follow this tutorial with more recent releases of LAMMPS–GUI and LAMMPS; older versions may require some (minor) adjustments. These tutorials will be periodically updated to ensure compatibility and benefit from new features in the latest stable version of LAMMPS.
For some tutorials, external tools are required for plotting and visualization, as the corresponding functionality in LAMMPS–GUI is limited. Suitable tools for plotting include Python with Pandas/Matplotlib [22, 23], XmGrace, Gnuplot, Microsoft Excel, or LibreOffice Calc. For visualization, suitable tools include VMD [24, 25] and OVITO [26, 27].
2.3 About LAMMPS–GUI
LAMMPS–GUI is a graphical text editor, enhanced for editing LAMMPS input files and linked to the LAMMPS library, allowing it to run LAMMPS directly. The text editor is similar to other graphical editors, such as Notepad or Gedit, but offers the following enhancements for running LAMMPS:
Wizard dialogs to set up these tutorials
Auto-completion of LAMMPS commands and options
Context-sensitive online help
Syntax highlighting for LAMMPS input files
Syntax-aware line indentation
Editor switches working directory to that of input file
Visualization using LAMMPS’ built-in renderer
Start and stop simulations via mouse or keyboard
Monitoring of simulation progress and parallelization
Dynamic capture of LAMMPS output in a text window
Automatic plotting of thermodynamic data during runs
Capture of “dump image” outputs for animations
Export of thermodynamic data for external plotting
Inspection of binary restart files
Appendix A contains basic instructions for installation and using LAMMPS–GUI with the tutorials presented here. A complete description of all LAMMPS–GUI features can be found in the LAMMPS manual [16].
3 Content and links
The tutorials described in this article can be accessed at lammpstutorials.github.io, where additional exercises with solutions are also provided. All files and inputs required to follow the tutorials are available from a dedicated GitHub organization account, github.com/lammpstutorials. These files can also be downloaded by clicking «Start LAMMPS Tutorial X» (where X = 1…8) from the «Tutorials» menu of LAMMPS–GUI.
In the following, all LAMMPS input or console commands are formatted with a colored background. Keyboard shortcuts and file names are formatted in monospace, and LAMMPS–GUI options and menus are displayed in «quoted monospace».
3.1 Tutorial 1: Lennard-Jones fluid
The objective of this tutorial is to perform simple MD simulations using LAMMPS. The system consists of a Lennard-Jones fluid composed of neutral particles with two different effective diameters, contained within a cubic box with periodic boundary conditions (Fig. 1). In this tutorial, basic MD simulations in the microcanonical (NVE) and canonical (NVT) ensembles are performed, and basic quantities are calculated, including the potential and kinetic energies.
3.1.1 My first input
To run a simulation using LAMMPS, you need to write an input script containing a series of commands for LAMMPS to execute, similar to Python or Bash scripts. For clarity, the input scripts for this tutorial will be divided into five categories, which will be filled out step by step. To set up this tutorial, select «Start LAMMPS Tutorial 1» from the «Tutorials» menu of LAMMPS–GUI, and follow the instructions. This will select (or create, if needed) a folder, place the initial input file initial.lmp in it, and open the file in the LAMMPS–GUI «Editor» window. It should display the following content:
Everything that appears after a hash symbol () is a comment and ignored by LAMMPS. LAMMPS–GUI will color such comments in red. These five categories are not required in every input script and do not necessarily need to be in that exact order. For instance, the Settings and the Monitoring categories could be inverted, or the Monitoring category could be omitted. However, note that LAMMPS reads input files from top to bottom and processes each command immediately. Therefore, the Initialization and System definition categories must appear at the top of the input, and the Run category must appear at the bottom. Also, the specifics of some commands can change after global settings are modified, so the order of commands in the input script is important.
Initialization
In the first section of the script, called Initialization, global parameters for the simulation are defined, such as units, boundary conditions (e.g., periodic or non-periodic), and atom types (e.g., uncharged point particles or extended spheres with a radius and angular velocities). These commands must be executed before creating the simulation box or they will cause an error. Similarly, many LAMMPS commands may only be entered after the simulation box is defined. Only a limited number of commands may be used in both cases. Update the initial.lmp file so that the Initialization section appears as follows:
Note 3.1.
Strictly speaking, none of the four commands specified in the Initialization section are mandatory, as they correspond to the default settings for their respective global properties. However, explicitly specifying these defaults is considered good practice to avoid confusion when sharing input files with other LAMMPS users.
The first line, units lj, specifies the use of reduced units, where all quantities are dimensionless. This unit system is a popular choice for simulations that explore general statistical mechanical principles, as it emphasizes relative differences between parameters rather than representing any specific material. The second line, dimension 3, specifies that the simulation is conducted in 3D space, as opposed to 2D, where atoms are confined to move only in the xy-plane. The third line, atom_style atomic, designates the atomic style for representing simple, individual point particles. In this style, each particle is treated as a point with a mass, making it the most basic atom style. Other atom styles can incorporate additional attributes for atoms, such as charges, bonds, or molecule IDs, depending on the requirements of the simulated model. The last line, boundary p p p, indicates that periodic boundary conditions are applied along all three directions of space, where the three p stand for , , and , respectively. Alternatives are fixed non-periodic (f), shrink-wrapped non-periodic (s), and shrink-wrapped non-periodic with minimum (m). For non-periodic boundaries, different options can be assigned to each dimension, making configurations like boundary p p fm valid for systems such as slab geometries.
Note 3.2.
Each LAMMPS command is accompanied by extensive online documentation that lists and discusses the different options for that command. Most LAMMPS commands also have default settings that are applied if no value is explicitly specified. The defaults for each command are listed at the bottom of its documentation page. From the LAMMPS–GUI editor buffer, you can access the documentation by right-clicking on a line containing a command (e.g., units lj) and selecting «View Documentation for ‘units’». This action should prompt your web browser to open the corresponding URL for the online manual. A screenshot of this context menu is shown in Fig. 2.
System definition
The next step is to create the simulation box and populate it with atoms. Modify the System definition category of initial.lmp as shown below:
The first line, region simbox (…), defines a region named simbox that is a block (i.e., a rectangular cuboid) extending from -20 to 20 units along all three spatial dimensions. The second line, create_box 2 simbox, initializes a simulation box based on the region simbox and reserves space for two types of atoms. In LAMMPS, every atom is assigned an atom type property. This property selects which force field parameters (here, the Lennard-Jones parameters, and ) are applied to each pair of atoms via the pair_coeff command (see below). We discuss in Tutorial 2 how this applies to many-body pair styles, and in Tutorial 3 how this applies to Coulomb interactions.
Note 3.3.
From this point on, the number of atom types is “locked in”, and any command referencing an atom type larger than 2 will trigger an error. While it is possible to allocate more atom types than needed, you must assign a mass and provide force field parameters for each atom type. Failing to do so will cause LAMMPS to terminate with an error.
The third line, create_atoms (…), generates 1500 atoms of type 1 at random positions within the simbox region. The integer 34134 is a seed for the internal random number generator, which can be changed to produce different sequences of random numbers and, consequently, different initial atom positions. The fourth line adds 100 atoms of type 2. Both create_atoms commands use the optional argument overlap 0.3, which enforces a minimum distance of 0.3 length units between the randomly placed atoms. This constraint helps avoid “close contacts” between atoms, which can lead to excessively large forces and simulation instability. Each created atom in LAMMPS is automatically assigned a unique atom ID, which serves as a numerical identifier to distinguish individual atoms throughout the simulation. Atom IDs by default have the range from 1 to the total number of atoms, but this is not enforced. Deleting atoms, for example, causes “holes” in the list of atom IDs.
Note 3.4.
Another way to define a system in LAMMPS, besides the create_atoms commands, is to import an existing topology file containing atomic coordinates as well as, optionally, other attributes such as atomic velocities and the force field parameters using the read_data command, as shown in Tutorial 2.
Settings
Next, we specify the settings for the two atom types. Modify the Settings category of initial.lmp as follows:
The two mass commands assign a mass of 1.0 and 5.0 units to the atoms of type 1 and 2, respectively. The third line, pair_style lj/cut 4.0, specifies that the atoms will be interacting through a Lennard-Jones (LJ) potential with a cut-off equal to length units [28, 29]:
where is the inter-particle distance, is the depth of the potential well that determines the interaction strength, and is the distance at which the potential energy equals zero. The indices and refer to pairs of atoms with the corresponding atom types. The fourth line, pair_coeff 1 1 1.0 1.0, specifies the Lennard-Jones coefficients for interactions between pairs of atoms that both have atom type 1: the energy parameter and the distance parameter . Similarly, the last line sets the Lennard-Jones coefficients for interactions between atoms of type 2, , and .
Note 3.5.
By default, LAMMPS calculates the mixed Lennard-Jones coefficients for pairs of atoms having distinct atom types using geometric averages: , . In the present case, , and . Other rules can be selected using the pair_modify command.
Single-point energy
The system is now fully parameterized. Let us complete the two remaining categories, Monitoring and Run, by adding the following lines to initial.lmp:
The thermo 10 command instructs LAMMPS to print thermodynamic information to the console every specified number of steps, in this case, every 10 simulation steps. The thermo_style custom command defines the specific outputs, which in this case are the step number (step), total energy (etotal), and pressure (press). The run 0 post no command instructs LAMMPS to initialize forces and energy without actually running the simulation. The post no option disables the post-run summary and statistics output.
Note 3.6.
The ‘thermodynamic information’ printed by LAMMPS using thermo_style custom keywords refers to instantaneous values of the specified thermodynamic properties at each output step, not cumulative averages. However, LAMMPS also allows to reference a wide variety of custom data from compute styles, fix styles, and variables. These can be used for on-the-fly analysis, including cumulative and sliding-window averages.
You can now run LAMMPS (see subsection A.3 for details on running LAMMPS). The simulation should finish quickly, and, with the default settings, LAMMPS–GUI will open two windows: one displaying the console output and another with a chart. The «Output» window will display information from the executed commands, including the total energy and pressure at step 0, as specified by the thermodynamic data request. Since no actual simulation steps were performed, the «Charts» window will be empty.
Snapshot Image
At this point, you can create a snapshot image of the current system using the «Image Viewer» window, which can be accessed by clicking the «Create Image» button in the «Run» menu. The image viewer works by instructing LAMMPS to render an image of the current system using its internal rendering library via the dump image command. The resulting image is then displayed, with various buttons available to adjust the view and rendering style. The image shown in Fig. 3 was created this way. This will always capture the current state of the system. Save the image for future comparisons by clicking the «Save as» button in the «File» menu.
Energy minimization
Now, replace the run 0 post no command line with the following minimize command:
This tells LAMMPS to perform an iterative energy minimization of the system. Specifically, LAMMPS will compute the forces on all atoms and then update their positions according to a selected algorithm using the computed forces, aiming to reduce the potential energy. By default, LAMMPS uses the conjugate gradient (CG) algorithm [30]. The simulation will stop as soon as one of the four minimizer criteria is met. LAMMPS will then report which stopping criterion was satisfied, along with selected system properties at both the initial and final steps. Note that, except for trivial systems, minimization algorithms will find a local minimum rather than the global minimum.
Run the minimization and observe that LAMMPS–GUI captures the output and update the chart in real time (see Fig. 4). This run executes quickly (depending on your computer’s capabilities) and thus LAMMPS–GUI may fail to capture some of the thermodynamic data. In that case, use the «Preferences» dialog to reduce the data update interval and switch to single-threaded, unaccelerated execution in the «Accelerators» tab. You can repeat the run; each new attempt will start fresh, resetting the system and re-executing the script from the beginning.
The potential energy, , decreases from a positive value to a negative value (Figs. 4 and 5 a). Note that during the energy minimization, the potential energy equals the total energy of the system, , since the kinetic energy, , is zero. The initially positive potential energy is expected, as the atoms are created at random positions within the simulation box, with some in very close proximity to each other. This proximity results in a large initial potential energy due to the repulsive branch of the Lennard-Jones potential [i.e., the term in Eq. (1)]. As the energy minimization progresses, the energy decreases - first rapidly - then more gradually, before plateauing at a negative value. This indicates that the atoms have moved to reasonable distances from one another.
Note 3.7.
Since the thermo_style command also includes the press keyword, you can switch from plotting the total energy to displaying the pressure by selecting «Press» in the «Select data» drop-down menu of the «Charts» window.
Create and save a snapshot image of the simulation state after the minimization, and compare it to the initial image. You should observe that the atoms are “clumping together” as they move toward positions of lower potential energy.
Molecular dynamics
After the energy minimization, any overlapping atoms are displaced, and the system is ready for a molecular dynamics simulation. To continue from the result of the minimization step, append the MD simulation commands to the same input script, initial.lmp. Add the following lines immediately after the minimize command:
Since LAMMPS reads inputs from top to bottom, these lines will be executed after the energy minimization. Therefore, there is no need to re-initialize or re-define the system. The thermo command is called a second time to update the output frequency from 10 to 50 as soon as PART B of the simulation starts. In addition, a new thermo_style command is introduced to specify the thermodynamic information LAMMPS should print during during PART B. This adjustment is made because, during molecular dynamics, the system exhibits a non-zero temperature (and consequently a non-zero kinetic energy , keyword ke), which are useful to monitor. The pe keyword represents the potential energy of the system, , such that .
Then, add a second Run category by including the following lines in PART B of initial.lmp:
The fix nve command updates the positions and velocities of the atoms in the group all at every step. More specifically, this command integrates Newton’s equations of motion using the velocity-Verlet algorithm. The group all is a default group that contains all atoms. The last two lines specify the value of the timestep and the number of steps for the run, respectively, for a total duration of 250 time units.
Note 3.8.
Since the only command affecting forces and velocities in the present script is fix nve, and periodic boundary conditions are applied in all directions, the MD simulation will be performed in the microcanonical (NVE) statistical mechanical ensemble, which maintains a constant number of particles and a fixed box volume. In this ensemble, the system does not exchange energy with anything outside the simulation box.
Run the simulation using LAMMPS. Initially, the system is not equilibrated, as the potential energy decreases while the kinetic energy increases. After approximately 40 000 steps, the values for both kinetic and potential energy plateau, indicating that the system has reached equilibrium, with the total energy fluctuating around a certain constant value.
Now, we change the second Run section to (note the smaller number of MD steps):
The new command adds a Langevin thermostat to the atoms in the group all, with a target temperature of 1.0 temperature units throughout the run (the two numbers represent the target temperature at the beginning and at the end of the run, which results in a temperature ramp if they differ) [31]. A damping parameter of 0.1 is used. It determines how rapidly the temperature is relaxed to its desired value. In a Langevin thermostat, the atoms are subject to friction and random noise (in the form of randomly added velocities). Since a constant friction term removes more kinetic energy from fast atoms and less from slow atoms, the system will eventually reach a dynamic equilibrium where the kinetic energy removed and added are about the same. The number 10917 is a seed used to initialize the random number generator used inside of fix langevin; you can change it to perform statistically independent simulations. In the presence of a thermostat, the MD simulation will be performed in the canonical or NVT ensemble.
Run the simulation again using LAMMPS–GUI. From the information printed in the «Output» window, one can see that the temperature starts from 0 but rapidly reaches the requested value and stabilizes itself near temperature units. One can also observe that the potential energy, , rapidly decreases during energy minimization (see also Fig. 5 a). After the molecular dynamics simulation starts, increases until it reaches a plateau value of about -0.25. The kinetic energy, , is equal to zero during energy minimization and then increases rapidly during molecular dynamics until it reaches a plateau value of about 1.5 (Fig. 5 d).
Note 3.9.
All simulations presented in these tutorials are deliberately kept short so they can be executed on a personal computer. These runs are not intended to produce statistically meaningful results, and should not be considered suitable for publication (see for instance Ref. 32).
Trajectory visualization
So far, the simulation has been mostly monitored through the analysis of thermodynamic information. To better follow the evolution of the system and visualize the trajectories of the atoms, let us use the dump image command to create snapshot images during the simulation. We have already explored the «Image Viewer» window. Open it again and adjust the visualization to your liking using the available buttons. Now you can copy the commands used to create this visualization to the clipboard by either using the «Ctrl-D» keyboard shortcut or selecting «Copy dump image command» from the «File» menu. This text can be pasted into the into the Monitoring section of PART B of the initial.lmp file. This may look like the following:
The ‘’ characters at the end are used to extend the commands across multiple lines. These two commands tell LAMMPS to generate NetPBM format images every 100 steps. The two type keywords are for color and diameter, respectively. Run the initial.lmp using LAMMPS again, and a new window named «Slide Show» will pop up. It will show each image created by the dump image as it is created. After the simulation is finished (or stopped), the slideshow viewer allows you to animate the trajectory by cycling through the images. The window also allows you to export the animation to a movie (provided the FFMpeg program is installed) and to bulk delete those image files.
The rendering of the system can be further adjusted using the many options of the dump image command. For instance, the value for the shiny keyword is used to adjust the shininess of the atoms, the box keyword adds or removes a representation of the box, and the view and zoom keywords adjust the camera (and so on).
3.1.2 Improving the script
Let us improve the input script and perform more advanced operations, such as specifying initial positions for the atoms and restarting the simulation from a previously saved configuration.
Control the initial atom positions
Open the improved.min.lmp, which was downloaded during the tutorial setup. This file contains the Part A of the initial.lmp file, but without any commands in the System definition section:
We want to create the atoms of types 1 and 2 in two separate regions. To achieve this, we need to add two region commands and then reintroduce the create_atoms commands, this time using the new regions instead of the simulation box region to place the atoms:
The side in and side out keywords are used to define regions representing the inside and outside of the cylinder of radius 10 length units, respectively. Then, append a sixth section titled Save system at the end of the file, ensuring that the write_data command is placed after the minimize command:
Note 3.10.
A key improvement to the input is the addition of the write_data command. This command writes the state of the system to a text file called improved.min.data. This .data file will be used later to restart the simulation from the final state of the energy minimization step, eliminating the need to repeat the system creation and minimization.
Run the improved.min.lmp file using LAMMPS–GUI. At the end of the simulation, a file called improved.min.data is created. You can view the contents of this file from LAMMPS–GUI, by right-clicking on the file name in the editor and selecting the entry «View file improved.min.data».
The created .data file contains all the information necessary to restart the simulation, such as the number of atoms, the box size, the masses, and the pair coefficients. This .data file also contains the final positions of the atoms, along with their IDs and types, within the Atoms section. The first five columns of the Atoms section correspond (from left to right) to the atom indexes (from 1 to the total number of atoms, 1150), the atom types (1 or 2 here), and the positions of the atoms , , . The last three columns are image flags that keep track of which atoms crossed the periodic boundary. The exact format of each line in the Atoms section depends on the choice of atom_style, which determines which per-atom data is set and stored internally in LAMMPS.
Note 3.11.
Instead of the write_data command, you can also use the write_restart command to save the state of the simulation to a binary restart file. Binary restart files are more compact, faster to write, and contain more information, making them often more convenient to use. For example, the choice of atom_style or pair_style is recorded, so those commands do not need to be issued before reading the restart. Note however that restart files are not expected to be portable across LAMMPS versions or platforms. Therefore, in these tutorials, and with the exception of Tutorial 3, we primarily use write_data to provide you with a reference copy of the data file that works regardless of your LAMMPS version and platform.
Restarting from a saved configuration
To continue a simulation from the saved configuration, open the improved.md.lmp file, which was downloaded during the tutorial setup. This file contains the Initialization part from initial.lmp and improved.min.lmp:
Since we read most of the information from the data file, we don’t need to repeat all the commands from the System definition and Settings categories. The exception is the pair_style command, which now must come before the simulation box is defined, meaning before the read_data command. Add the following lines to improved.md.lmp:
By visualizing the system (see Fig. 6), you may have noticed that some atoms left their original region during minimization. To start the simulation from a clean slate, with only atoms of type 2 inside the cylinder and atoms of type 1 outside the cylinder, let us delete the misplaced atoms by adding the following commands to the System definition section of the improved.md.lmp:
The first two region commands recreate the previously defined regions, which is necessary since regions are not saved by the write_data command. The first two group commands create groups containing all the atoms of type 1 and all the atoms of type 2, respectively. The next two group commands create atom groups based on their positions at the beginning of the simulation, i.e., when the commands are being read by LAMMPS. The last two group commands create atom groups based on the intersection between the previously defined groups. Finally, the two delete_atoms commands delete the atoms of type 1 located inside the cylinder and the atoms of type 2 located outside the cylinder, respectively.
Since LAMMPS has a limited number of custom groups (30), it is good practice to delete groups that are no longer needed. This can be done by adding the following four commands to improved.md.lmp:
Let us monitor the number of atoms of each type inside the cylinder as a function of time by creating the following equal-style variables:
The equal-style variables are expressions evaluated during the run and return a number. Here, they are defined to count the number of atoms of a specific group within the cyl_in region.
Note 3.12.
The n1_in and n2_in defined above are equal-style variables, which evaluate a numerical expression using the count() function. Other LAMMPS variable styles include atom, index, file, loop, string, and vector.
In addition to counting the atoms in each region, we will also extract the coordination number of type 2 atoms around type 1 atoms. The coordination number measures the number of type 2 atoms near type 1 atoms, defined by a cutoff distance. Taking the average provides as a good indicator of the degree of mixing in a binary mixture. This is done using two compute commands: the first counts the coordinated atoms, and the second calculates the average over all type 1 atoms. Add the following lines to improved.md.lmp:
The compute reduce ave command is used to average the per-atom coordination number calculated by the compute coord/atom command. Compute commands do not print or output anything by themselves, nor are they automatically executed; they require a “consumer” command that references the compute. In this case, the first compute is referenced by the second, and we reference the second in a thermo_style custom command (see below).
Note 3.13.
LAMMPS compute commands can produce a wide variety of data and one can identify the category from the name of the compute style: global data (no suffix), local data (/local suffix), per-atom data (/atom suffix), per-chunk data (/chunk suffix), per-grid data (/grid suffix). In the example above, the compute coord/atom produces per-atom data, which is used as input for compute reduce which returns global data. For global data three kinds of data exists: scalars (single values), vectors (one-dimensional arrays), or arrays (two-dimensional tables). When referencing results of a compute, you can use indices: for example, c_mycompute refers to the entire scalar, vector, or array, and c_mycompute[1] refers to its first element or column (in case of vector or array). In some cases also wildcards like “*” can be used to, for instance, refer to all elements of a vector instead of having specify all elements individually. In general, “consumer” commands (fix styles or dump styles, variables, or other compute styles) can only work with certain data types or need to have keywords set to select which data to use. You need to check the documentation of each command to ensure compatibility.
There is no need for a Settings section, as the settings are taken from the .data file. Finally, let us complete the script by adding the following lines to improved.md.lmp:
The two variables n1_in, n2_in, along with the compute sumcoor12, were added to the list of information printed during the simulation. Additionally, images of the system will be created with slightly less saturated colors than the default ones.
Finally, add the following lines to improved.md.lmp:
Here, there are a few more differences from the previous simulation. First, the velocity create command assigns an initial velocity to each atom. The initial velocity is chosen so that the average initial temperature is equal to 1.0 temperature units. The additional keywords ensure that no linear momentum (mom yes) is given to the system and that the generated velocities are distributed according to a Gaussian distribution. Another improvement is the zero yes keyword in the Langevin thermostat, which ensures that the total random force applied to the atoms is equal to zero. These steps are important to prevent the system from starting to drift or move as a whole.
Note 3.14.
A bulk system with periodic boundary conditions is expected to remain in place. Accordingly, when computing the temperature from the kinetic energy, we use degrees of freedom since there is no global translation. In a drifting system, some of the kinetic energy is due to the drift, which means the system itself cools down. In extreme cases, the system can freeze while its center of mass drifts very quickly. This phenomenon is sometimes referred to as the “flying ice cube syndrome” [7].
Run improved.md.lmp and observe the mixing of the two populations over time (see also Fig. 7). From the variables n1_in and n2_in, you can track the number of atoms in each region as a function of time (Fig. 8 a). To view their evolution, select the entries «v_n1_in» or «v_n2_in» in the «Data» drop-down menu in the «Charts» window of LAMMPS–GUI.
In addition, as the mixing progresses, the average coordination number between atoms of types 1 and 2 increases from about to (Fig. 8 b). This indicates that, over time, more and more particles of type 1 come into contact with particles of type 2, as expected during mixing. This can be observed using the entry «c_sumcoor12» in the «Charts» drop-down menu.
Experiments
Here are some suggestions for further experiments with this system that may lead to additional insights into how different systems are configured and how various features function:
Use a Nosé-Hoover thermostat (fix nvt) instead of a Langevin thermostat (fix nve + fix langevin).
Omit the energy minimization step before starting the MD simulation using either the Nosé-Hoover or the Langevin thermostat.
Apply a thermostat to only one type of atoms each and observe the temperature for each type separately.
Append an NVE run (i.e. without any thermostat) and observe the energy levels.
Note 3.15.
In contrast to the fix nve command, which integrates Newton’s equations of motion without any thermostatting, the fix nvt command adds a Nosé-Hoover thermostat to control the system temperature.
Another useful experiment is coloring the atoms in the «Slide Show» according to an observable, such as their respective coordination numbers. To do this, replace the dump and dump_modify commands with the following lines:
Run LAMMPS again. Atoms of type 1 are now colored based on the value of c_coor12, which is mapped continuously from turquoise to yellow and red for atoms with the highest coordination (Fig. 9). In the definition of the variable v_coor12, atoms of type 2 are all assigned a value of -1, and will therefore always be colored their default blue.
3.2 Tutorial 2: Pulling on a carbon nanotube
In this tutorial, the system of interest is a small, single-walled carbon nanotube (CNT) in an empty box (Fig. 10). The CNT is strained by imposing a constant velocity on the edge atoms. To illustrate the difference between conventional and reactive force fields, this tutorial is divided into two parts: in the first part, a conventional molecular force field (called OPLS-AA [10]) is used and the functional form of the bonded potential ensures that the bonds between the atoms of the CNT are unbreakable. In the second part, a reactive, many-body force field (called AIREBO [11]) is used, which allows chemical bonds to break under large strain.
To set up this tutorial, select «Start Tutorial 2» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. This will select a folder, create one if necessary, and place several files into it. The initial input file, set up for a single-point energy calculation, will also be loaded into the editor under the name unbreakable.lmp. Additional files are a data file containing the CNT topology and geometry, named unbreakable.data, a parameters file named unbreakable.inc, as well as the scripts required for the second part of the tutorial.
3.2.1 Unbreakable bonds
With most conventional molecular force fields, the chemical bonds between atoms are defined at the start of the simulation and remain fixed, regardless of the forces applied to the atoms. In this tutorial, these bonds are explicitly specified in the .data file, which is read using the read_data command (see below). Bonds are typically modeled as springs following Hooke’s law with equilibrium distances , force constants , and bond potential energy . Additionally, angular and dihedral interactions are often imposed to preserve the molecular structure by maintaining the relative orientations of neighboring atoms.
The LAMMPS input
After completing the setup, the editor should display the following content:
The chosen unit system is real (therefore distances are in Ångströms (Å), times in femtoseconds (fs), and energies in (kcal/mol)), the atom_style is molecular (therefore atoms are point particles that can form bonds with each other), and the boundary conditions are fixed. The boundary conditions do not matter here, as the box boundaries were placed far from the CNT. Just like in the previous tutorial, Lennard-Jones fluid, the pair style is lj/cut (i.e. a Lennard-Jones potential with cutoff) and its cutoff is set to 14 Å, which means that only the atoms closer than this distance interact through the Lennard-Jones potential.
The bond_style, angle_style, dihedral_style, and improper_style commands specify the different potentials used to constrain the relative positions of the atoms. The special_bonds command sets the weighting factors for the Lennard-Jones interactions between atoms sitting one, two, or three bonds away from each other, respectively. This is done for convenience when parameterizing the force constants for bonds, angles, and so on. By excluding the non-bonded (Lennard-Jones) interactions for these pairs, those interactions do not need to be considered when determining the force constants.
The read_data command imports the unbreakable.data file that should have been downloaded next to unbreakable.lmp during the tutorial setup. This file contains information about the box size, atom positions, as well as the identity of the atoms that are linked by bonds, angles, dihedrals, and impropers interactions. It was created using VMD and TopoTools [33].
Note 3.16.
Bonds, angles, dihedrals, and impropers in LAMMPS are assigned types and IDs, just like atoms. The ID uniquely identifies each interaction instance, while the type determines which parameters (from the bond_coeff, angle_coeff, etc. commands) are applied. In this tutorial, these types and IDs are specified in the .data file and read by the read_data command.
Note 3.17.
The format details of the different sections in a data file change with different settings. In particular, the Atoms section may have a different number of columns, or the columns may represent different properties when the atom_style is changed. To help users, LAMMPS and tools like VMD and TopoTools will add a comment (here # molecular) to the Atoms header line in the data files that indicates the intended atom_style. LAMMPS will print a warning when the chosen atom style does not match what is written in that comment.
The .data file does not contain any sections with potential parameters; thus, we need to specify the parameters of both the bonded and non-bonded potentials. The parameters we use are taken from the OPLS-AA (Optimized Potentials for Liquid Simulations-All-Atom) force field [10], and are given in a separate unbreakable.inc file (also downloaded during the tutorial setup). This file - that must be placed next to unbreakable.lmp - contains the following lines:
The pair_coeff command sets the parameters for non-bonded Lennard-Jones interactions between atoms type 1 to and . The bond_coeff provides the equilibrium distance and the spring constant for the harmonic potential imposed between two bonded carbon atoms. The potential is given by . The angle_coeff gives the equilibrium angle and constant for the potential between atoms forming an angle: . The dihedral_coeff and improper_coeff define the potentials for the constraints between 4 atoms.
Note 3.18.
Rather than copying the contents of the file into the input, we incorporate it using the include command. Using include allows us to conveniently reuse the parameter settings in other inputs or switch them with others. This will become more general when using type labels [12], which is shown in the next tutorial.
Prepare the initial state
In this tutorial, a deformation will be applied to the CNT by displacing the atoms located at its edges. To achieve this, we will first isolate the atoms at the two edges and place them into groups named rtop and rbot. Add the following lines to unbreakable.lmp, just before the run 0 command:
The first command includes all the atoms of type 1 (i.e. all the atoms here) in a group named carbon_atoms. The variable corresponds to the coordinate of the last atoms along minus 0.5 Å, and to the coordinate of the first atoms along plus 0.5 Å. Then, three regions are defined, corresponding to the following: , (rbot, for region bottom), (rmid, for region middle), and (rtop, for region top).
Note 3.19.
So far, variables have been referenced dynamically during the run using the v_ prefix, which evaluates the variable as it evolves over time. Here, a dollar sign ($) is used to expand the variable immediately at the time the input script is read.
Finally, let us define 3 groups of atoms corresponding to the atoms in each of the 3 regions by adding to unbreakable.lmp just before the run 0 command:
With the three set commands, we assign unique, otherwise unused molecule IDs to atoms in those three groups. A molecule ID is an integer that groups atoms into a ‘molecule’ for bookkeeping purposes, and can be useful for tracking and post-processing. We will use these IDs later to assign different colors to these groups of atoms.
Run the simulation using LAMMPS. The number of atoms in each group is given in the «Output» window. It is an important check to make sure that the number of atoms in each group corresponds to what is expected, as shown here:
Finally, to start from a less ideal state and create a system with some defects, let us randomly delete a small fraction of the carbon atoms. To avoid deleting atoms that are too close to the edges, let us define a new region named rdel that starts at from the CNT edges:
The delete_atoms command randomly deletes of the atoms from the rdel group, here about 10 atoms (compare the top and the middle panels in Fig. 11).
The molecular dynamics
Let us give an initial temperature to the atoms of the group cnt_mid by adding the following commands to unbreakable.lmp:
Re-setting the atom IDs is necessary before using the velocity command when atoms were deleted, which is done here with the reset_atoms command. The velocity command assigns random initial velocities to the atoms of the middle group cnt_mid from a uniform distribution, ensuring an initial temperature of for these atoms.
Let us specify the thermalization and the dynamics of the system. Add the following lines into unbreakable.lmp:
The fix nve commands are applied to the atoms of cnt_top and cnt_bot, respectively, and will ensure that the positions of the atoms from these groups are recalculated at every step. The fix nvt does the same for the cnt_mid group, while also applying a Nosé-Hoover thermostat with desired temperature of 300 K [34, 35].
Note 3.20.
The Nosé-Hoover thermostat only controls the temperature of the atoms belonging to the specified cnt_mid group. Atoms outside this group are not affected.
To immobilize the atoms at the edges, let us add the following commands to unbreakable.lmp:
The two setforce commands cancel the forces applied on the atoms of the two edges, respectively. The cancellation of the forces is done at every step, and along all 3 directions of space, , , and , due to the use of 0 0 0. Although the force on these atoms is set to zero, the fix stores the force vector acting on the group before cancellation, which can later be extracted for analysis (see below). The two velocity commands set the initial velocities along , , and to 0 for the atoms of cnt_top and cnt_bot, respectively. As a consequence of these last four commands, the atoms of the edges will remain immobile during the simulation (or at least they would if no other command was applied to them).
Note 3.21.
The velocity set command adjusts the velocities of a group of atoms immediately but has no effect during the simulation. When velocity set is used in combination with setforce 0 0 0, as is the case here, the initial velocity will persist during the entire simulation, thus producing a constant velocity motion or no motion at all.
Outputs
Next, to measure the strain and stress applied to the CNT, let us create a variable for the distance between the two edges, as well as a variable for the force applied on the edges:
Here, the force is extracted from the fixes mysf1 and mysf2 using f_, similarly to the use of v_ to call a variable, and c_ to call a compute, as seen in Tutorial 1.
Let us also add a dump image command to visualize the system every 500 steps:
Let us run a small equilibration step to bring the system to the required temperature before applying any deformation. Replace the run 0 post no command in unbreakable.lmp with the following lines:
With the thermo_modify command, we specify to LAMMPS that the temperature mid of the middle group, cnt_mid, must be outputted, instead of the temperature of the entire system. This choice is motivated by the presence of frozen parts with an effective temperature of 0 K, which makes the average temperature of the entire system less relevant. The thermo_modify command also imposes the use of the YAML format that can easily be read by Python (see below).
Let us impose a constant velocity deformation on the CNT by combining the velocity set command with previously defined fix setforce. Add the following lines in the unbreakable.lmp file, right after the last run 5000 command:
The chosen velocity for the deformation is , or .
Run the simulation using LAMMPS. As can be seen from the variable , the length of the CNT increases linearly over time for (Fig. 12 a), as expected from the imposed constant velocity. What you observe in the «Slide Show» windows should resembles Fig. 11. The total energy of the system shows a non-linear increase with once the deformation starts, which is expected from the typical dependency of bond energy with bond distance, (Fig. 12 b).
Importing YAML log file into Python
Let us import the simulation data into Python, and generate a stress-strain curve. Here, the stress is defined as , where is the surface area of the CNT, and Å the CNT radius. The strain is defined as , where is the initial CNT length.
Right-click inside the «Output» window, and select «Export YAML data to file». Call the output unbreakable.yaml, and save it within the same folder as the input files, where a Python script named unbreakable-yaml-reader.py should also be located. When executed using Python, this .py file first imports the unbreakable.yaml file. Then, a certain pattern is identified and stored as a string character named ‘docs’. The string is then converted into a list, and and are extracted. The stress and strain are then calculated, and the result is saved in a data file named unbreakable.dat using the NumPy ‘savetxt’ function. ‘thermo[0]’ can be used to access the information from the first minimization run, and ‘thermo[1]’ to access the information from the second MD run. The data extracted from the unbreakable.yaml file can then be used to plot the stress-strain curve, see Fig. 13.
3.2.2 Breakable bonds
When using a conventional molecular force field, as we have just done, the bonds between the atoms are non-breakable. Let us perform a similar simulation and deform a small CNT again, but this time with a reactive force field that allows bonds to break if the applied deformation is large enough.
Input file initialization
Open the input named breakable.lmp that should have been downloaded next to unbreakable.lmp during the tutorial setup. There are only a few differences with the previous input. First, the AIREBO force field requires the metal units setting instead of real for OPLS-AA. A second difference is the use of atom_style atomic instead of molecular, since no explicit bond information is required with AIREBO. The following commands are setting up the AIREBO force field:
Note 3.22.
The AIREBO force field is a many-body potential, where interactions are not only between pairs of atoms, but also triples and quadruples representing angle and dihedral interactions. This means that there are different rules for the pair_coeff command: there must be only one command that covers all permutations of atom types by using two ’*’ wildcards. After the potential file follows a list of elements. These element names are used to look up the parameter sets in the potential file. There must be a list with as many elements as atom types following the filename. In our system, there is only one atom type (1), which is mapped to the element ’C’ in the pair_coeff command. Which elements are supported is determined by the contents of the potential file.
Here, CH.airebo is the file containing the parameters for AIREBO, and must be placed next to breakable.lmp.
Note 3.23.
With metal units, time values are in units of picoseconds ( s) instead of femtoseconds ( s) in the case of real units. It is important to keep this in mind when setting parameters that are expressed units containing time, such as the timestep or the time constant of a thermostat, or velocities.
Since bonds, angles, and dihedrals do not need to be explicitly set when using AIREBO, some simplification must be made to the .data file. The new .data file is named breakable.data, and must be placed within the same folder as the input file. Just like unbreakable.data, the breakable.data contains the information required for placing the atoms in the box, but no bond/angle/dihedral information. Another difference between the unbreakable.data and breakable.data files is that, here, a larger distance of 120 Å was used for the box size along the -axis, to allow for larger deformation of the CNT.
Start the simulation
Here, let us perform a similar deformation as the previous one. In breakable.lmp, replace the run 0 post no line with:
Note the relatively small timestep of ps ( fs) used. Reactive force fields like AIREBO usually require a smaller timestep than conventional ones. When running breakable.lmp with LAMMPS, you can see that the temperature deviates from the target temperature of at the start of the equilibration, but that after a few steps, it reaches the target value.
Note 3.24.
Bonds cannot be displayed by the dump image when using the atom_style atomic, as it contains no bonds. A tip for displaying bonds with the present system using LAMMPS is provided at the end of the tutorial. You can also use external tools like VMD or OVITO (see the tip for tutorial 3).
Launch the deformation
After equilibration, let us set the velocity of the edges equal to (or ) and run for a longer duration than previously. Add the following lines into breakable.lmp:
Run the simulation. Some bonds are expected to break before the end of the simulation (Fig. 14).
Looking at the evolution of the energy, one can see that the total energy is initially increasing with the deformation. When bonds break, the energy relaxes abruptly, as can be seen near in Fig. 15 a. Using a similar script as previously, i.e., unbreakable-yaml-reader.py, import the data into Python and generate the stress-strain curve (Fig. 15 b). The stress-strain curve reveals a linear (elastic) regime where for , and a non-linear (plastic) regime for .
Tip: bonds representation with AIREBO
In the input file solution/breakable-with-tip.lmp, which is an alternate solution for breakable.lmp, a trick is used to represent bonds while using AIREBO. A detailed explanation of the script is beyond the scope of the present tutorial. In short, the trick is to use AIREBO with the molecular atom style, and use the fix bond/break and fix bond/create/angle commands to update the status of the bonds during the simulation:
This “hack” works because AIREBO does not pay any attention to bonded interactions and computes the bond topology dynamically inside the pair style. Thus adding bonds of bond style zero does not add any interactions but allows the visualization of them with dump image. It is required to change the special_bonds setting to disable any neighbor list exclusions as they are common for force fields with explicit bonds.
3.3 Tutorial 3: Polymer in water
The goal of this tutorial is to use LAMMPS to solvate a small hydrophilic polymer molecule (PEG - polyethylene glycol) in a reservoir of water (Fig. 16). Once the water reservoir is properly equilibrated at the desired temperature and pressure, the polymer molecule is added and a constant stretching force is applied to both ends of the polymer. The evolution of the polymer length is measured as a function of time. The GROMOS 54A7 force field [36] is used for the PEG, the SPC/Fw model [37] is used for the water, and the long-range Coulomb interactions are solved using the PPPM solver [38]. This tutorial was inspired by a publication by Liese and coworkers, in which molecular dynamics simulations are compared with force spectroscopy experiments, see Ref. 39.
Note 3.25.
When mixing different force fields, as is done here with GROMOS and SPC/Fw, users should exercise caution. The choices made in these tutorials prioritize progressive learning of LAMMPS functionality over strict physical accuracy. While GROMOS is commonly used with water models from the SPC family [40], the inter-compatibility of force fields is not generally guaranteed.
3.3.1 Preparing the water reservoir
In this tutorial, the water reservoir is first prepared in the absence of the polymer. A rectangular box of water is created and equilibrated at ambient temperature and pressure. The SPC/Fw water model is used [37], which is a flexible variant of the rigid SPC (simple point charge) model [41]. To set up this tutorial, select «Start Tutorial 3» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to water.lmp:
With the unit style real, masses are in g/mol, distances in Å, time in fs, and energies in kcal/mol. With the atom_style full, each atom is a dot with a mass and a charge that can be linked by bonds, angles, dihedrals, and/or impropers. The bond_style, angle_style, and dihedral_style commands define the potentials for the bonds, angles, and dihedrals used in the simulation, here harmonic. With the pair_style named lj/cut/coul/long, atoms interact through both a Lennard-Jones (LJ) potential and Coulomb interactions. The value of is the cutoff, and the kspace_style command defines the long-range solver for the Coulomb interactions [42]. Finally, the special_bonds command, which was already seen in Tutorial 2, sets the LJ and Coulomb weighting factors for the interaction between neighboring atoms.
Note 3.26.
With Coulomb interactions, additional rules apply to the pair_coeff command: (a) atom type values only matter for assignment of LJ potential parameters; (b) for Coulomb interactions, there are no parameters outside the cutoff, and when using a coul/long pair style, that cutoff can only be set globally for all atoms with the pair_style command; (c) for Coulomb interactions, only the per-atom charge and any special_bonds exclusions are relevant.
Let us create a 3D simulation box of dimensions , and make space for 8 atom types (2 for the water, 6 for the polymer), 7 bond types (1 for the water, 6 for the polymer), 8 angle types (1 for the water, 7 for the polymer), and 4 dihedral types (only for the polymer). Copy the following lines into water.lmp:
The extra/x/per/atom commands reserve memory for adding bond topology data later. We use the file parameters.inc to set all the parameters (masses, interaction energies, bond equilibrium distances, etc). Thus add to water.lmp the line:
Note 3.27.
This tutorial uses type labels [12] to map each numeric atom type to a string (see the parameters.inc file): labelmap atom 1 OE 2 C 3 HC 4 H 5 CPos 6 OAlc 7 OW 8 HW Therefore, the oxygen and hydrogen atoms of water (respectively types 7 and 8) can be referred to as ‘OW’ and ‘HW’, respectively. Similar maps are used for the bond types, angle types, and dihedral types.
Let us create water molecules. To do so, let us import a molecule template called water.mol and then randomly create 700 molecules. Add the following lines into water.lmp:
The first parameter is 0, meaning that the atom types from the water.mol file will be used. The overlap 1.0 option of the create_atoms command ensures that no atoms are placed exactly in the same position, as this would cause the simulation to crash. The maxtry 50 asks LAMMPS to try at most 50 times to insert the molecules, which is useful in case some insertion attempts are rejected due to overlap. In some cases, depending on the system and the values of overlap and maxtry, LAMMPS may not create the desired number of molecules. Always check the number of created atoms in the log file (or in the «Output» window), where you should see:
When LAMMPS fails to create the desired number of molecules, a WARNING appears. The molecule template called water.mol must be downloaded and saved next to water.lmp. This template contains the necessary structural information of a water molecule, such as the number of atoms, or the IDs of the atoms that are connected by bonds and angles.
Then, let us organize the atoms of types OW and HW of the water molecules in a group named H2O and perform a small energy minimization. The energy minimization is mandatory here because of the small overlap value of 1 Å chosen in the create_atoms command. Add the following lines into water.lmp:
Resetting the step of the simulation to 0 using the reset_timestep command is optional. It is used here because the number of iterations performed by the minimize command is usually not a round number, since the minimization stops when one of four criteria is reached, which can disrupt the intended frequency of outputs such as dump commands that depend on the timestep count. We will use fix npt to control the temperature and pressure of the molecules with a Nosé-Hoover thermostat and barostat, respectively [34, 35, 43]. Add the following line into water.lmp:
The fix npt allows us to impose both a temperature of (with a damping constant of ), and a pressure of 1 atmosphere (with a damping constant of ). With the iso keyword, the three dimensions of the box will be re-scaled isotropically, maintaining the same proportion in all directions.
Let us output the system into images by adding the following commands to water.lmp:
Let us also extract the volume and density, among others, every 500 steps:
With the real units system, the volume is in Å3, and the density is in g/cm3.
Finally, let us set the timestep to 1.0 fs, and run the simulation for 15 ps by adding the following lines into water.lmp:
The final state is saved in a binary file named water.restart. Run the input using LAMMPS. The system reaches its equilibrium temperature after just a few picoseconds, and its equilibrium density after approximately 10 picoseconds (Fig. 17). A snapshot of the equilibrated system can also be seen in Fig. 18.
Note 3.28.
The binary file created by the write_restart command contains the complete state of the simulation, including atomic positions, velocities, and box dimensions (similar to write_data), but also the groups, the compute, or the atom_style. Use the «Inspect Restart» option of the LAMMPS–GUI to vizualize the content saved in water.restart.
3.3.2 Solvating the PEG in water
Now that the water reservoir is equilibrated, we can safely add the PEG polymer to the water. The PEG molecule topology was downloaded from the ATB repository [44, 40]. It has a formula , and the parameters are taken from the GROMOS 54A7 force field [36] (Fig. 19).
Open the file named merge.lmp that was downloaded alongside water.lmp during the tutorial setup. It only contain one line:
Most of the commands that were initially present in water.lmp, such as the units of the atom_style commands do not need to be repeated, as they were saved within the .restart file. There is also no need to re-include the parameters from the .inc file. The kspace_style command, however, is not saved by the write_restart command and must be repeated. Since Ewald summation is not the most efficient choice for such dense system, let us use PPPM (for particle-particle particle-mesh) for the rest of the tutorial. Add the following command to merge.lmp:
Using the molecule template for the polymer called peg.mol, let us create a single molecule in the middle of the box by adding the following commands to merge.lmp:
Let us create a group for the atoms of the PEG (the previously created group H2O was saved by the restart and can be omitted):
Water molecules that are overlapping with the PEG must be deleted to avoid future crashing. Add the following line into merge.lmp:
Here the value of 2.0 Å for the overlap cutoff was fixed arbitrarily and can be chosen through trial and error. If the cutoff is too small, the simulation will crash because atoms that are too close to each other undergo forces that can be extremely large. If the cutoff is too large, too many water molecules will unnecessarily be deleted.
Let us use the fix npt to control the temperature, as well as the pressure by allowing the box size to be rescaled along the -axis:
Let us also use the recenter command to always keep the PEG at the position :
Note 3.29.
Note that the recenter command has no impact on the dynamics, it simply repositions the frame of reference so that any drift of the system is ignored, which can be convenient for visualizing and analyzing the system. However, be aware that using fix recenter can sometimes mask underlying issues in the simulation, such as a net momentum or the so-called “flying ice cube syndrome” [7].
Let us create images of the systems:
Finally, to perform a short equilibration and save the final state to a .restart file, add the following lines to the input:
Run the simulation using LAMMPS. From the outputs, you can make sure that the temperature remains close to the target value of throughout the entire simulation, and that the volume and total energy are almost constant, indicating that the system was in a reasonable configuration from the start. See a snapshot of the system in Fig. 20.
3.3.3 Stretching the PEG molecule
Here, a constant force is applied to both ends of the PEG molecule until it stretches. Open the file named pull.lmp, which only contains two lines:
Next, we’ll create new atom groups, each containing a single oxygen atom. The atoms of type OAlc correspond to the hydroxyl (alcohol) group oxygen atoms located at the ends of the PEG molecule, which we will use to apply the force. Add the following lines to pull.lmp:
These lines identify the oxygen atoms (type OAlc) at the ends of the PEG molecule and calculates their center of mass along the -axis. It then divides these atoms into two groups, end1 (i.e., the OAlc atom to the right of the center) and end2 (i.e., the OAlc atom to the right of the center), for applying force during the stretching process.
Add the following dump command to create images of the system:
Let us use a single Nosé-Hoover thermostat applied to all the atoms, and let us keep the PEG in the center of the box, by adding the following lines to pull.lmp:
To investigate the stretching of the PEG molecule, let us compute its radius of gyration [45] and the angles of its dihedral constraints using the following commands:
The radius of gyration can be directly printed with the thermo_style command:
By contrast with the radius of gyration (compute rgyr), the dihedral angle (compute dphi) is returned as a vector by the compute dihedral/local command and must be written to a file using the dump local command.
Finally, let us simulate 15 picoseconds without any external force:
This initial run will serve as a benchmark to quantify the changes caused by the applied force in later steps. Next, let us apply a force to the two selected oxygen atoms using two addforce commands, and then run the simulation for an extra 15 ps:
Each applied force has a magnitude of , corresponding to . This value was chosen to be sufficiently large to overcome both the thermal agitation and the entropic contributions from the molecules.
Run the pull.lmp file using LAMMPS. From the generated images of the system, you should observe that the PEG molecule eventually aligns in the direction of the applied force (as seen in Fig. 21). The evolutions of the radius of gyration over time indicates that the PEG quickly adjusts to the external force (Fig. 22 a). Additionally, from the values of the dihedral angles printed in the pull.dat file, you can create a histogram of dihedral angles for a specific type. For example, the angle for dihedrals of type 1 (C-C-OE-C) is shown in Fig. 22 b.
Tip: using external visualization tools
Trajectories can be visualized using external tools such as VMD or OVITO [25, 27]. To do so, the IDs and positions of the atoms must be regularly written to a file during the simulation. This can be accomplished by adding a dump command to the input file. For instance, create a duplicate of pull.lmp and name it pull-with-tip.lmp. Then, replace the existing dump and dump_modify commands with:
Running the pull-with-tip.lmp file using LAMMPS will generate a trajectory file named pull.lammpstrj, which can be opened in OVITO or VMD.
Note 3.30.
Since the default trajectory dump file does not contain information about topology and elements, it is usually preferred to first write out a data file and import it directly (in the case of OVITO) or convert it to a PSF file (for VMD). This allows the topology to be loaded before adding the trajectory file to it. When using LAMMPS–GUI, this process can be automated through the «View in OVITO» or «View in VMD» options in the «Run» menu. Afterwards only the trajectory dump needs to be added. Alternatively, the dump custom command can be combined with dump command to include element names in the dump file and simplify visualization.
Note 3.31.
Microstates collected during a simulation in the form of a trajectory can be analyzed within LAMMPS using the rerun command. This is particularly useful, for example, for computing properties not set up in the original simulation without having to run it again. A possible use of the rerun command is estimating the self-diffusion coefficient by using the compute msd command [1].
3.4 Tutorial 4: Nanosheared electrolyte
The objective of this tutorial is to simulate an electrolyte nanoconfined and sheared between two walls (Fig. 23). The density and velocity profiles of the fluid in the direction normal to the walls are extracted to highlight the effect of confining a fluid on its local properties. This tutorial demonstrates key concepts of combining a fluid and a solid in the same simulation. A major difference from the previous tutorial, Polymer in water, is that here a rigid four-point water model named TIP4P/2005 is used [13].
Note 3.32.
Four-point water models such as TIP4P/2005 are widely used as they offer a good compromise between accuracy and computational cost [46].
3.4.1 System preparation
The fluid and walls must first be generated, followed by equilibration at the desired temperature and pressure.
System generation
To set up this tutorial, select «Start Tutorial 4» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to create.lmp:
These lines are used to define the most basic parameters, including the atom style, the forms of the non-bonded, bond, and angle potentials, as well as other specifics of the non-bonded interactions. Here, lj/cut/tip4p/long imposes a Lennard-Jones potential with a cut-off at and a long-range Coulomb potential. The parameters O, H, O-H, and H-O-H correspond respectively to the oxygens, hydrogens, O-H bonds, and H-O-H angle constraints of the water molecules; their definitions, provided by the labelmap commands, will be clarified below.
So far, the commands are relatively similar to those in the previous tutorial, Polymer in water, with two major differences: the use of lj/cut/tip4p/long instead of lj/cut/coul/long, and pppm/tip4p instead of pppm. When using lj/cut/tip4p/long and pppm/tip4p, the interactions resemble the conventional Lennard-Jones and Coulomb interactions, except that they are specifically designed for the four-point water model. As a result, LAMMPS automatically adds the fourth point to the water molecules, assigning type O atoms as oxygen and type H atoms as hydrogen. The fourth massless atom (M) of the TIP4P water molecule does not have to be defined explicitly, and the value of corresponds to the O-M distance of the TIP4P-2005 water model [13]. All other atoms in the simulation are treated as usual, with long-range Coulomb interactions. Another novelty, here, is the use of kspace_modify slab 3.0 that is combined with the non-periodic boundaries along the coordinate: boundary p p f. With the slab option, the system is treated as periodical along , but with an empty volume inserted between the periodic images of the slab, and the interactions along effectively turned off.
Let us create the box and the label maps by adding the following lines to create.lmp:
The lattice command defines the unit cell. Here, the face-centered cubic (fcc) lattice with a scale factor of 4.04 has been chosen for the future positioning of the atoms of the walls. The region command defines a geometric region of space. By choosing and , and because we have previously chosen a lattice with a scale factor of 4.04, the region box extends from to along the direction. The create_box command creates a simulation box with 5 types of atoms: the oxygen and hydrogen of the water molecules, the two ions (, ), and the atoms from the walls. The simulation contains 1 type of bond and 1 type of angle (both required by the water molecules). The parameters for these bond and angle constraints will be given later. The extra/ (…) keywords are for memory allocation. Finally, the labelmap commands assign alphanumeric type labels to each numeric atom type, bond type, and angle type, concepts already introduced in previous tutorials.
Now, we can add atoms to the system. First, let us create two sub-regions corresponding respectively to the two solid walls, and create a larger region from the union of the two regions. Then, let us create atoms of type WALL within the two regions. Add the following lines to create.lmp:
Atoms will be placed in the positions of the previously defined lattice, thus forming fcc solids.
To add the water molecules, the molecule template called water.mol must be located next to create.lmp. The template contains all the necessary information concerning the water molecule, such as atom positions, bonds, and angles. Add the following lines to create.lmp:
Within the last three lines, a region named rliquid is created based on the last defined lattice, fcc 4.04. rliquid will be used for introducing the water molecules in the simulation domain. The molecule command opens up the molecule template called water.mol, and names the associated molecule h2omol. The new molecules are placed on the fcc 4.04 lattice by the create_atoms command. The first parameter is 0, meaning that the atom types from the water.mol file will be used. The number 482793 is a seed that is required by LAMMPS, it can be any positive integer.
Finally, let us create 30 ions (15 and 15 ) in between the water molecules, by adding the following commands to create.lmp:
Each create_atoms command will add 15 ions at random positions within the rliquid region, ensuring that there is no overlap with existing molecules. Feel free to increase or decrease the salt concentration by changing the number of desired ions. To keep the system charge neutral, always insert the same number of and , unless there are other charges in the system. The charges of the newly added ions are specified by the two set commands.
Before starting the simulation, we need to define the parameters of the simulation: the mass of the 5 atom types (O, H, , , and wall), the pairwise interaction parameters (in this case, for the Lennard-Jones potential), and the bond and angle parameters. Copy the following lines into create.lmp:
Both parameters.inc and groups.inc files must be located next to create.lmp.
The parameters.inc file contains the masses, as follows:
Each mass command assigns a mass in g/mol to an atom type. The parameters.inc file also contains the pair coefficients:
Each pair_coeff assigns the depth of the LJ potential (in kcal/mol), and the distance (in Ångströms) at which the particle-particle potential energy is 0. As noted in previous tutorials, with the important exception of pair_coeff O WALL, pairwise interactions were only assigned between atoms of identical types. By default, LAMMPS calculates the pair coefficients for the interactions between atoms of different types (i and j) by using geometric average: , . However, if the default value of was used for , the solid walls would be extremely hydrophilic, causing the water molecules to form dense layers. As a comparison, the water-water energy is only . Therefore, to make the walls less hydrophilic, the value of was reduced.
Finally, the parameters.inc file contains the following two lines:
The bond_coeff command, used here for the O-H bond of the water molecule, sets both the spring constant of the harmonic potential and the equilibrium bond distance of . The force constant can be 0 for a rigid water molecule because the SHAKE algorithm, which will be used in the input at a later step, will constrain the intramolecular structure of the water molecules (see below) [47, 48]. Similarly, the angle_coeff command for the H-O-H angle of the water molecule sets the force constant of the angular harmonic potential to 0 and the equilibrium angle to .
Alongside parameters.inc, the groups.inc file contains several group commands to define groups of atoms based on their types:
The groups.inc file also defines the walltop and wallbot groups, which contain the WALL atoms located in the and regions, respectively:
Currently, the fluid density between the two walls is slightly too high. To avoid excessive pressure, let us add the following lines into create.lmp to delete about of the water molecules:
To create an image of the system, add the following dump image into create.lmp (see also Fig. 24):
Finally, add the following lines into create.lmp:
The run 0 command initializes the simulation, which is required for cleanly saving the state, but it does not advance positions or velocities. The write_data command generates a file called system.data containing the information required to restart the simulation from the final configuration produced by this input file. With the nocoeff option, the parameters from the force field are not included in the .data file. Run the create.lmp file using LAMMPS, and a file named create.data will be created alongside create.lmp.
Energy minimization
Let us move the atoms and place them in more energetically favorable positions before starting the actual molecular dynamics simulation. Open the equilibrate.lmp file that was downloaded alongside create.lmp during the tutorial setup. Same as before, it contains the following lines:
The only difference from the previous input is that, instead of creating a new box and new atoms, we open the previously created create.data file.
Now, let us use the SHAKE algorithm to maintain the shape of the water molecules [47, 48] by adding the following line to the script.
Here the SHAKE algorithm applies to the O-H bond and the H-O-H angle of the water molecules. The kbond keyword specifies the force constant that will be used to apply a restraint force when used during minimization. This last keyword is important here, because the spring constants of the rigid water molecules were set to 0 (see the parameters.inc file).
Note 3.33.
LAMMPS provides several ways to keep molecules rigid during a simulation. The fix shake command is appropriate and efficient for constraining individual bonds or bonds and angles within small molecules like water while using a per-atom time integration fix command like fix nve or fix nvt. However, it fails for linear molecules like CO2 or constraining larger or more complex objects. In such cases, the fix rigid family of commands can be used to perform time integration for translation and rotation of groups of atoms as rigid bodies.
Let us also create images of the system and control the printing of thermodynamic outputs by adding the following lines to equilibrate.lmp:
Let us perform an energy minimization by adding the following lines to equilibrate.lmp:
When running the equilibrate.lmp file with LAMMPS, you should observe that the total energy of the system is initially very high but rapidly decreases. From the generated images of the system, you will notice that the atoms and molecules are moving to adopt more favorable positions.
System equilibration
Let us equilibrate further the entire system by letting both fluid and wall relax at ambient temperature. Here, the commands are written within the same equilibrate.lmp file, right after the reset_timestep command.
Let us do a molecular dynamics simulation using the Nosé-Hoover thermostat. Add the following lines to equilibrate.lmp:
As mentioned previously, the fix recenter does not influence the dynamics, but will keep the system in the center of the box, which makes the visualization easier. Then, add the following lines into equilibrate.lmp for the trajectory visualization:
The undump command is used to cancel the previous dump command. Then, a new dump command with a larger dumping period is used.
Note 3.34.
Just like the undump command can cancel an active dump, other objects defined in a LAMMPS input script can be cancelled when no longer needed. For example, you can use unfix to remove a previously defined fix, and uncompute to delete a compute.
To monitor the system equilibration, let us print the distance between the two walls. Add the following lines to equilibrate.lmp:
The first two variables extract the z coordinate of the centers of mass of the two walls. The deltaz variable is then used to calculate the difference between the two variables walltopz and wallbotz, i.e. the distance in the z direction between the two centers of mass of the walls.
Finally, let us run the simulation for 30 ps by adding a run command to equilibrate.lmp:
Run the equilibrate.lmp file using LAMMPS. Both the pressure and the distance between the two walls show oscillations at the start of the simulation but eventually stabilize at their equilibrium values toward the end of the simulation (Fig. 25).
Note 3.35.
Note that it is generally recommended to run a longer equilibration. In this case, the slowest process in the system is likely ionic diffusion. Therefore, the equilibration period should, in principle, exceed the time required for the ions to diffuse across the size of the pore, i.e. . Using as the final pore size and as the typical diffusion coefficient for sodium chloride in water at room temperature [49], one finds that the equilibration should be on the order of one nanosecond.
3.4.2 Imposed shearing
From the equilibrated configuration, let us impose a lateral motion on the two walls and shear the electrolyte. Open the last input file named shearing.lmp. It starts as follows:
To address the dynamics of the system, add the following lines to shearing.lmp:
One key difference with the previous input is that, here, two thermostats are used, one for the fluid (mynvt1) and one for the solid (mynvt2). The combination of fix_modify with compute temp ensures that the correct temperature values are used by the thermostats. Using compute commands for the temperature with temp/partial 0 1 1 is intended to exclude the coordinate from the thermalization, which is important since a large velocity will be imposed along the direction.
Then, let us impose the velocity of the two walls by adding the following commands to shearing.lmp:
The setforce commands cancel the forces on walltop and wallbot in the direction. As a result, the atoms in these two groups will not experience any forces along from their interaction with rest of the system. Consequently, in the absence of external forces, these atoms will conserve the initial velocities imposed by the two velocity commands. As seen previously, although the forces on these atoms are set to zero, the fix setforce still stores the forces acting on the group before cancellation, which can later be extracted for analysis (see below).
Finally, let us generate images of the systems and print the values of the forces exerted by the fluid on the walls, as given by f_mysf1[1] and f_mysf2[1]. Add these lines to shearing.lmp:
Let us also extract the density and velocity profiles using the chunk/atom and ave/chunk commands. When deployed as below, these commands discretize the simulation domain into spatial bins and compute and output average properties of the atoms belonging to each bin, here the velocity along (vx) within the bins. Add the following lines to shearing.lmp:
Here, a bin size of is used for the density profiles generated by the ave/chunk commands, and three .dat files are created for the water, the walls, and the ions, respectively. With values of 10 15000 200000, the velocity vx will be evaluated every 10 steps during the final 150,000 steps of the simulations. The result will be averaged and printed only once at the 200,000th step.
Run the simulation using LAMMPS. The averaged velocity profile for the fluid is plotted in Fig. 26. As expected for such a Couette flow geometry, the fluid velocity increases linearly along , and is equal to the walls velocities at the fluid-solid interfaces (no-slip boundary conditions).
From the force applied by the fluid on the solid, one can extract the stress within the fluid, which enables the measurement of its viscosity according to
where is the stress applied by the fluid on the shearing wall, and the shear rate [50]. Here, the shear rate is approximately (Fig. 26), the average force on each wall is given by f_mysf1[1] and f_mysf2[1] and is approximately . Using a surface area for the walls of , one obtains an estimate for the shear viscosity for the confined fluid of using Eq. (2).
Note 3.36.
The viscosity calculated at such a high shear rate may differ from the expected bulk value. In general, it is recommended to use a lower value for the shear rate. Note that for lower shear rates, the signal-to-noise ratio is smaller, and longer simulations are needed. Another point to consider is that the viscosity of a fluid next to a solid surface is typically larger than in bulk due to interaction with the walls [51]. Therefore, one expects the present simulation to yield a viscosity that is slightly higher than what would be measured in the absence of walls.
3.5 Tutorial 5: Reactive silicon dioxide
The objective of this tutorial is to demonstrate how the reactive force field ReaxFF can be used to calculate the partial charges of a system undergoing deformation, as well as the formation and breaking of chemical bonds [14, 52]. The system simulated in this tutorial is a block of silicon dioxide (Fig. 27) which is deformed until it ruptures. Particular attention is given to the evolution of atomic charges during deformation, with a focus on tracking chemical reactions resulting from the deformation over time.
3.5.1 Prepare and relax
The first step is to relax the structure with ReaxFF, which which will be achieved using molecular dynamics. To ensure the system equilibrates properly, we will monitor certain parameters over time, such as the system volume. To set up this tutorial, select «Start Tutorial 5» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to relax.lmp:
So far, the input is very similar to what was seen in the previous tutorials. Some basic parameters are defined (units and atom_style), and a .data file is imported by the read_data command.
The initial topology given by silica.data corresponds to a small amorphous silica structure. This structure was generated in a prior simulation using the Vashishta force field [53]. If you open the silica.data file, you will find in the Atoms section that all silicon atoms have a charge of , and all oxygen atoms have a charge of .
Note 3.37.
Assigning the same charge to all atoms of the same type is common with many force fields, including the force fields used in the previous tutorials. This changes once ReaxFF is used: the charge of each atom will adjust to its local environment through charge equilibration.
Next, copy the following three crucial lines into the relax.lmp file:
In this case, the pair_style reaxff is used without a control file (see note below). The safezone and mincap keywords are added to prevent allocation issues, which sometimes can trigger segmentation faults and bondchk errors. The pair_coeff command uses the ffield.reax.CHOFe file, which should have been downloaded during the tutorial set up. Finally, the fix qeq/reaxff is used to perform charge equilibration [54], which occurs at every step. The values 0.0 and 10.0 represent the low and the high cutoffs, respectively, and is the tolerance, i.e., the precision to which the atomic charges are equilibrated during the charge equilibration process. The maxiter sets an upper limit to the number of attempts to equilibrate the charge.
Note 3.38.
The pair_style reaxff command optionally accepts a control file, which defines control variables such as global parameters of the ReaxFF potential, as well as performance and output settings. If no control file is provided, as in this tutorial, LAMMPS uses its default values, which correspond to those in Adri van Duin’s original stand-alone ReaxFF code [14].
Next, add the following commands to the relax.lmp file to track the evolution of the charges during the simulation:
The definition of the equal style variables qSi and qO make use of functions pre-defined within LAMMPS that allow calculating the total charge of atoms belonging to a group (charge()) and the total number of atoms in the group (count()). To print the averaged charges qSi and qO using the thermo_style command, and create images of the system. Add the following lines to relax.lmp:
Here, the atoms are colored by their charges q, ranging from royal blue (when ) to orange-red (when ).
We can generate histograms of the charges for each atom type using fix ave/histo commands:
The fix ave/histo command samples values over a group of atoms and builds a histogram over a specified range divided into bins. In this tutorial, it is used to monitor the charge distributions of silicon and oxygen atoms. The parameters 10 500 5000 specify how often the histogram is updated and averaged, -1.5 2.5 set the value range, 1000 is the number of bins, and v_vq is the variable being histogrammed.
We can also use fix reaxff/species to evaluate what species are present within the simulation. It will be useful later when the system is deformed, and bonds are broken:
Here, the information will be printed every 5 steps in a file called relax.species. Let us perform a very short run using the anisotropic NPT command and relax the density of the system:
The write_data command is used with the nofix keyword to print a data file without extra sections from the reaxff/species command. Run the relax.lmp file using LAMMPS. As seen from relax.species, only one species is detected, called O384Si192, representing the entire system.
Note 3.39.
With the aniso keyword, the three dimensions of the simulation box can change independently. This is particularly relevant for solids and other systems where anisotropic stresses may develop.
As the simulation progresses, the charge of every atom fluctuates because it is adjusting to the local environment of the atom (Fig. 29 a). It is also observed that the averaged charges for silicon and oxygen atoms fluctuate significantly at the beginning of the simulation, corresponding to a rapid change in the system volume, which causes interatomic distances to shift quickly (Fig. 29 b). The atoms with the most extreme charges are located at structural defects, such as dangling oxygen groups (Fig. 28). Finally, the generated .histo files can be used to plot the probability distributions, (see Fig. 30 a).
3.5.2 Deform the structure
Let us apply a deformation to the structure to force some bonds to break (and eventually re-assemble). Open the deform.lmp file, which must contain the following lines:
The only difference with the previous relax.lmp file is the path to the relax.data file.
Next, let us use fix nvt instead of fix npt to apply a Nosé-Hoover thermostat without a barostat:
Here, no barostat is used because the change in the box volume will be imposed by the fix deform, see below.
Let us run for 5000 steps without deformation, then apply the fix deform to progressively elongate the box along the -axis during 25000 steps. Add the following line to deform.lmp:
The fix deform command applies a continuous deformation by elongating the simulation box along the x-axis at a constant engineering shear strain rate, specified by erate, of .
Run the deform.lmp file using LAMMPS. During the deformation, the charge values progressively evolve until the structure eventually breaks down. After the structure breaks down, the charges equilibrate near new average values that differ from the initial averages (Fig. 31 a). The difference between the initial and the final charges can be explained by the presence of defects, as well as new solid/vacuum interfaces, and the fact that surface atoms typically have different charges compared to bulk atoms (Fig. 32). You can also see a sharp increase in temperature during the rupture of the material (Fig. 31 b).
You can examine the charge distribution after deformation, as well as during deformation (Fig. 30 b). As expected, the final charge distribution slightly differs from the previously calculated one. If no new species were formed during the simulation, the deform.species file should look like this:
Sometimes, molecules are formed during the deformation. If this occurs, a new column O2 appears in the deform.species file.
3.5.3 Decorate the surface
Under ambient conditions, some of the surface atoms become chemically passivated by forming covalent bonds with hydrogen (H) atoms [55]. We will add hydrogen atoms randomly to the cracked silica and observe how the system evolves. To do so, we first need to modify the previously generated data file deform.data and make space for a third atom type. Copy deform.data, name the copy deform-mod.data, and modify the first lines of deform-mod.data as follows:
Open the decorate.lmp file, which must contain the following lines:
The displace_atoms command is used to move the center of the crack near the center of the box. This step is optional but makes for a nicer visualization. A different value for the shift may be needed in your case, depending on the location of the crack. A difference with the previous input is that three atom types are specified in the pair_coeff command, i.e. Si O H.
Then, let us adapt some familiar commands to measure the charges of all three types of atoms, and output the charge values into log files:
The commands above are, once again, similar to the ones of the previous script. Here, the was added to the denominator of the variable qH to avoid dividing by 0 at the beginning of the simulation, as no hydrogen atoms exists in the simulation domain yet. Finally, let us create a loop with 10 steps, and create two hydrogen atoms at random locations at every step:
Run the simulation with LAMMPS. When the simulation is over, it can be seen from the decorate.species file that all the created hydrogen atoms reacted with the structure to form surface groups (such as hydroxyl (-OH) groups).
At the end of the simulation, hydroxyl (-OH) groups can be seen at the interfaces (Fig. 33).
3.6 Tutorial 6: Water adsorption in silica
The objective of this tutorial is to combine molecular dynamics and grand canonical Monte Carlo simulations to compute the adsorption of water molecules in cracked silica material (Fig. 34). This tutorial illustrates the use of the grand canonical ensemble in molecular simulation, an open ensemble where the number of atoms or molecules in the simulation box can vary. By using this combination, we simulate water in a nanoporous SiO2 structure at a specified chemical potential.
3.6.1 Generation of the silica block
To begin this tutorial, select «Start Tutorial 6» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to generate.lmp:
The main difference from some of the previous tutorials is the use of the Vashishta pair style. The Vashishta potential implicitly models atomic bonds through energy terms dependent on interatomic distances and angles [53].
Let us create a box for two atom types, Si of mass 28.0855 g/mol and O of mass 15.9994 g/mol. Add the following lines to generate.lmp:
In line with what is done in previous tutorials, the create_atoms commands are used to place 240 Si atoms and 480 O atoms, respectively, in the region previously defined. This corresponds to an initial density of approximately g/cm3, which is close to the expected final density of amorphous silica at 300 K.
Now, specify the potential parameters by indicating that the first atom type is Si and the second is O:
Ensure that the SiO.1990.vashishta file is located in the same directory as generate.lmp.
Next, add a dump image command to generate.lmp to follow the evolution of the system with time:
Let us also print the box volume and system density, alongside the temperature and total energy:
Finally, let us implement the annealing procedure which consists of three consecutive runs. This procedure was inspired by Ref. [56]. First, to melt the system, a run at is performed:
Next, a second run, during which the system is cooled down from to , is implemented as follows:
In this case, the initial and final target temperatures set for the Nosé-Hoover thermostat is different, causing it to evolve linearly within the number of timesteps evoked in the run command. In the third run, the system is equilibrated at the final desired conditions, and , using an anisotropic pressure coupling:
Here, an anisotropic barostat is used. As previously mentioned, anisotropic barostats adjust the dimensions independently, which is generally suitable for a solid phase.
Run the simulation using LAMMPS. From the «Charts» window, the temperature evolution can be observed, showing that it closely follows the desired annealing procedure (Fig. 35 a). The evolution of the box dimensions over time confirms that the box is deforming during the last stage of the simulation (Fig. 35 b). After the simulation completes, the final microstate attained during the dynamics and the system topology will be written to a LAMMPS data file called generate.data which will be located next to generate.lmp (Fig. 36).
3.6.2 Cracking the silica
Open the cracking.lmp file, which must contain the following familiar lines:
Let us progressively increase the size of the box in the direction, forcing the silica to deform and eventually crack. To achive this, the fix deform command is used, with a rate of . Add the following lines to the cracking.lmp file:
As discussed, the fix nvt command integrates the Nosé-Hoover equations of motion and is employed to control the temperature of the system. As observed from the generated images, the atoms progressively adjust to the changing box dimensions. At some point, bonds begin to break, leading to the appearance of dislocations (Fig. 37).
Note 3.40.
Although the Nosé-Hoover equations were originally formulated to sample the NVT ensemble, using the fix nvt command does not guarantee that a simulation actually samples the NVT ensemble.
3.6.3 Adding water
To add the water molecules to the silica, we will employ the Monte Carlo method in the grand canonical ensemble (GCMC). In short, the system is placed into contact with a virtual reservoir containing pure water at a given thermodynamic state, and multiple attempts to insert water molecules at random positions are made. In the grand canonical ensemble, each attempt is either accepted or rejected based on internal energy and chemical potential, considerations. For further details, please refer to classical textbooks like Ref. 1.
Adapting the pair style
For this next step, we need to specify the force field used to model the interactions in the system. The TIP4P/2005 model is employed for the water [13], while no interaction within silica is defined, as it will be seen farther below. This is because the atoms of the silica will remain frozen during this part of the simulation. Only the cross-interactions between water and silica need to be defined. Open the gcmc.lmp file, which should contain the following lines:
The PPPM solver [38] is specified with the kspace command, and is used to compute the long-range Coulomb interactions associated with tip4p/long. Finally, the form of the bond and angle potentials of the water molecules are defined; however, as previously discussed, these specifications are not critical since TIP4P/2005 is a rigid water model.
Note 3.41.
In practice, it is possible to use both vashishta and lj/cut/tip4p/long pair styles at the same time by employing the pair_style hybrid command. However, hybridizing force fields should be done with caution, as there is no guarantee that the resulting force field will produce meaningful results.
The water molecule template called H2O.mol must be downloaded and located next to gcmc.lmp.
Before going further, we need to make a few changes to our data file. Currently, the cracking.data file includes only two atom types, but we require four. Copy the previously generated cracking.data, and name the duplicate cracking-mod.data. Make the following changes to the beginning of cracking-mod.data to ensure it matches the following format (with 4 atom types, 1 bond type, 1 angle type, the proper type labels, and four masses):
Doing so, we anticipate that there will be 4 atom types in the simulations, with the oxygens and hydrogens of having types OW and HW, respectively. There will also be 1 bond type (OW-HW) and 1 angle type (OW-HW-HW). The extra bond, extra angle, and extra special lines are here for memory allocation.
We can now proceed to complete the gcmc.lmp file by adding the system definition:
After reading the data file and defining the h2omol molecule from the H2O.txt file, the create_atoms command is used to include three water molecules in the system. Then, add the following pair_coeff (and bond_coeff and angle_coeff) commands to gcmc.lmp in order to set the potential parameters:
Pair coefficients for the lj/cut/tip4p/long pair style are defined between O() and between H() atoms, as well as between O()-O() and Si()-O(). Thus, the fluid-fluid and the fluid-solid interactions will be adressed with by the lj/cut/tip4p/long potential. The bond_coeff and angle_coeff commands set the OW-HW bond length to 0.9572 Å, and the HW-OW-HW angle to , respectively [13].
Note 3.42.
The pair coefficients for interactions between Si() and O() are set by the first command, pair_coeff * * 0 0, which effectively means that they do not interact. This is acceptable here because the silica atoms remain frozen during this part of the tutorial.
Add the following lines to gcmc.lmp as well:
The number of oxygen atoms from water molecules (i.e. the number of molecules) is calculated by the nO variable. As already discussed in other tutorials, the SHAKE algorithm is used to maintain the shape of the water molecules over time [47, 48].
Note 3.43.
Here, a variable of style ‘atom’ is used. Such variable defines a per-atom property, i.e., it evaluates the specified expression separately for each atom. This is often used to select atoms based on their properties or types.
Finally, let us create images of the system using dump image:
GCMC simulation
To prepare for the GCMC simulation, let us add the following lines into gcmc.lmp:
Here, the fix nvt applies only to the water molecules, so the atoms in the silica remain fixed. The compute_modify command with the dynamic/dof yes option is used for water to account for the fact that the number of molecules is not constant.
Finally, let us use the fix gcmc and perform the grand canonical Monte Carlo steps. Add the following lines into gcmc.lmp:
The fix gcmc command performs grand canonical Monte Carlo moves to insert, delete, or swap molecules. Here, 100 attempts are made every 100 steps. The mol h2omol keyword specifies the molecule type being inserted/deleted, while shake shak enforces rigid molecular constraints during these moves. With the pressure 100 keyword, a fictitious reservoir with a pressure of 100 atmospheres is used. The tfac_insert option ensures the correct estimate for the temperature of the inserted water molecules by taking into account the internal degrees of freedom.
Note 3.44.
At a pressure of , the chemical potential of water vapor at can be calculated using as where is the standard chemical potential (typically taken at a pressure ), is the gas constant, is the temperature.
Finally, let us print some information and run for 25 ps:
The f_ keywords extract the Monte Carlo move statistics which is computed (and can be extracted) by the fix gcmc command.
Note 3.45.
When using the pressure argument, LAMMPS ignores the value of the chemical potential (here , which corresponds roughly to ambient conditions, i.e. to a relative humidity [57].) The large pressure value of 100 bars was chosen to ensure that some successful insertions of molecules would occur during the short duration of this simulation.
Running this simulation using LAMMPS, one can see that after a few GCMC steps, the number of molecules starts increasing. Once the crack is filled with water molecules, the total number of molecules reaches a plateau (Figs. 38-39). The final number of molecules depends on the imposed pressure, temperature, and the interaction between water and silica (i.e. its hydrophilicity). Note that GCMC simulations of such dense phases are usually slow to converge due to the very low probability of successfully inserting a molecule. Here, the short simulation duration was made possible by the use of a high pressure.
3.7 Tutorial 7: Free energy calculation
The objective of this tutorial is to measure the free energy profile of particles through a barrier potential using two methods: free sampling and umbrella sampling [58, 2, 1] (Fig. 40). To simplify the process and minimize computation time, the barrier potential will be imposed on the atoms using an additional force, mimicking the presence of a repulsive area in the middle of the simulation box without needing to simulate additional atoms. The procedure is valid for more complex systems and can be adapted to many other situations, such as measuring adsorption barriers near an interface or calculating translocation barriers through a membrane [59, 60, 61, 62, 63].
3.7.1 Method 1: Free sampling
The most direct way to estimate a free energy profile is to sample the Boltzmann distribution using a classical (i.e. unbiased) molecular dynamics simulation, and compute relative Gibbs free energies from the relative probabilities of states using
where is the free energy difference, is the gas constant, is the temperature, is the density, and is a reference density. As an illustration, let us apply this method to a simple system that consists of a particles in a box in the presence of a position-dependent repulsive force that makes the center of the box a less favorable area to explore.
Basic LAMMPS parameters
To begin this tutorial, select «Start Tutorial 7» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to free-sampling.lmp:
Here, we begin by defining variables for the Lennard-Jones parameters ( and ) and for the repulsive potential parameters , which are 0, , and [see Eqs. (4-5) below]. The cut-off value of was chosen to create a Weeks-Chandler-Andersen (WCA) potential, which is a truncated and purely repulsive LJ potential [64]. The potential is also shifted to be equal to 0 at the cut-off using the pair_modify command. The unit system is real, in which energy is in kcal/mol, distance in Ångströms, or time in femtosecond, has been chosen for practical reasons: the WHAM algorithm used in the second part of the tutorial automatically assumes the energy to be in kcal/mol.
Note 3.46.
The syntax $(...), where a dollar sign is followed by parentheses, allows you to evaluate a numeric formula immediately, without having to assign it to a named variable first.
System creation and settings
Let us define the simulation box and randomly add atoms by addying the following lines to free-sampling.lmp:
Note 3.47.
In the pair_coeff command, the first two asterisks * * indicate that the parameters apply to all atom types in the simulation.
The variables 0, , and , defined in the previous subsection, are used here to create the repulsive potential, restricting the atoms from exploring the center of the box:
Taking the derivative of the potential with respect to , we obtain the expression for the force that will be imposed on the atoms:
Fig. 41 shows the potential and force along the -axis. With , 0 is of the same order of magnitude as the thermal energy , where is the Boltzmann constant and is the temperature used in this simulation. Under these conditions, particles are expected to frequently overcome the energy barrier due to thermal agitation.
We impose the force to the atoms in the simulation using the fix addforce command. Add the following lines to free-sampling.lmp:
Next, we use the Newtonian equations of motion with a Langevin thermostat by combining the fix nve with a fix langevin command:
When combining these two commands, the MD simulation operates in the NVT ensemble, maintaining a constant number of atoms , constant volume , and a temperature that fluctuates around a target value.
Note 3.48.
LAMMPS documentation suggests using damping constants for thermostats that are approximately 100 times the timestep value. In this case, a value of 500 is used, resulting in a relatively weak coupling to the thermostat.
To ensure that the equilibration time is sufficient, we will track the evolution of the number of atoms in the central - energetically unfavorable - region, defined below under the name mymes, using the n_center variable:
A dump image command was also added for system visualization. The other commands should also be familiar from previous tutorials.
Finally, let us perform an equilibration of 50000 steps, using a timestep of , corresponding to a total duration of :
Run the simulation with LAMMPS. The number of atoms in the central region, , reaches its equilibrium value after approximately (Fig. 42). A snapshot of the equilibrated system is shown in Fig. 43.
Run and data acquisition
Once the system is equilibrated, we will record the density profile of the atoms along the -axis using the ave/chunk command. Add the following line to free-sampling.lmp:
Here, the chunk/atom command discretizes the simulation domain into spatial bins of size 2 Å along the direction, and the fix ave/chunk command outputs the number density of atoms within each bin to the file free-sampling.dat. The step count is reset to 0 using reset_timestep to synchronize it with the output times of fix ave/chunk. Run the simulation using LAMMPS.
Data analysis
Once the simulation is complete, the density profile from free-sampling.dat shows that the density in the center of the box is about two orders of magnitude lower than inside the reservoir (Fig. 44 a). Next, we plot , where is the the density ratio, and compare it with the imposed potential from Eq. (4) (Fig. 44 b). The reference density, , was estimated by measuring the density of the reservoir from the density profiles. The agreement between the MD results and the imposed energy profile is excellent, despite some noise in the central part, where fewer data points are available due to the repulsive potential.
The limits of free sampling
Increasing the value of 0 reduces the average number of atoms in the central region, making it difficult to achieve a high-resolution free energy profile within reasonable simulation times. For example, running the same simulation with , corresponding to , results in no atoms exploring the central part of the simulation box during the simulation. In such a case, employing an enhanced sampling method is recommended, as done in the next section.
3.7.2 Method 2: Umbrella sampling
Umbrella sampling is a biased molecular dynamics method in which additional forces are added to a chosen atom to force it to explore the more unfavorable areas of the system [58, 2, 1]. Here, to encourage one of the atoms to explore the central region of the box, we apply a potential and force it to move along the -axis. The chosen path is called the axis of reaction. Several simulations (called windows) will be conducted with varying positions for the center of the applied biasing. The results will be analyzed using the weighted histogram analysis method (WHAM) [65, 66], which allows for the removal of the biasing effect and ultimately deduces the unbiased free energy profile.
LAMMPS input script
Open the file named umbrella-sampling.lmp, which should contain the following lines:
The first difference from the previous case is the larger value for the repulsive potential parameter 0, which makes the central area of the system very unlikely to be visited by free particles. The second difference is the introduction of the variable , which will be used for the biasing potential.
Let us create a simulation box with two atom types, including a single particle of type 2, by adding the following lines to umbrella-sampling.lmp:
Next, we assign the same mass and LJ parameters to both atom types 1 and 2, and place the atoms of type 2 into a group named topull:
Then, the same potential and force are applied to all the atoms, together with the same fix nve and fix langevin commands:
Next, we perform a brief equilibration to prepare for the umbrella sampling run:
So far, our code resembles that of Method 1, except for the additional particle of type 2. Particles of types 1 and 2 are identical. However, the particle of type 2 will also be exposed to the biasing potential , which forces it to explore the central part of the box (Fig. 45), thus justifying the definition of two atom types.
Now, we create a loop with 15 steps and progressively move the center of the bias potential by increments of 0.4 nm. Add the following lines to umbrella-sampling.lmp:
The definition of a variable of loop style serves the same purpose as in (Tutorial 5), and we highlight here the particular utility of using its value to distinguish the files written by the fix ave_time command for the different bias potentials. The spring command imposes the additional harmonic potential with the previously defined spring constant to the atoms in the group topull. The center of the harmonic potential, , successively takes values from to . For each value of , an equilibration step of 40 ps is performed, followed by a step of 400 ps during which the position of the particle of type 2 along the -axis, , is saved in data files named umbrella-sampling.i.dat, where ranges from 1 to 15. Run the umbrella-sampling.lmp file using LAMMPS.
Note 3.49.
The value of should be chosen with care: if is too small the particle won’t follow the biasing potential, and if is too large there will be no overlapping between the different windows, leading to poor reconstruction of the free energy profile.
WHAM algorithm
To generate the free energy profile from the particle positions saved in the umbrella-sampling.i.dat files, we use the WHAM [65, 66] algorithm as implemented by Alan Grossfield [67]. You can download it from Alan Grossfield’s website. Make sure you download the WHAM code version 2.1.0 or later which introduces the units command-line option used below. The executable called wham generated by following the instructions from the website must be placed next to umbrella-sampling.lmp. To apply the WHAM algorithm to our simulation, we need a metadata file containing:
the paths to all the data files,
the values of ,
the values of .
Download the umbrella-sampling.meta file and save it next to umbrella-sampling.lmp. Then, run the WHAM algorithm by typing the following command in the terminal:
where -30 and 30 are the boundaries, 50 is the number of bins, 1e-8 is the tolerance, and 119.8 is the temperature in Kelvin. A file called umbrella-sampling.dat is created, containing the free energy profile in kcal/mol. The resulting PMF can be compared with the imposed potential , showing excellent agreement (Fig. 46). Remarkably, this excellent agreement is achieved despite the very short calculation time and the high value for the energy barrier. Achieving similar results through free sampling would require performing extremely long and computationally expensive simulations.
3.8 Tutorial 8: Reactive Molecular Dynamics
The goal of this tutorial is to create a model of a carbon nanotube (CNT) embedded in a polymer melt made of polystyrene (PS) (Fig. 47). The REACTER protocol is used to simulate the polymerization of styrene monomers, and the polymerization reaction is followed in time [68, 15, 69]. In contrast with AIREBO (Tutorial 2) and ReaxFF (Tutorial 5), the REACTER protocol relies on the use of a classical force field that does not inherently model bond formation or breaking, but instead couples with an external algorithm to simulate polymerization reactions.
3.8.1 Creating the system
To begin this tutorial, select «Start Tutorial 8» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to mixing.lmp:
The class2 pair_styles compute a 6/9 Lennard-Jones potential [70]. The class2 bond, angle, dihedral, and improper styles are used as well, see the documentation for a description of the respective potential form they, each, prescribe. The tail yes option adds long-range van der Waals tail corrections to the energy and pressure. The mix sixthpower imposes the following mixing rule for the calculation of the cross coefficients:
Let us read the CNT.data file, which contains a periodic single-walled CNT. Add the following line to mixing.lmp:
The CNT is approximately in diameter and in length, oriented along the -axis. The simulation box is initially 12.0 nm in the two other dimensions before densification, making it straightforward to fill the box with styrene. To add 200 styrene molecules to the simulation box, we will use the styrene.mol molecule template file. Include the following commands to mixing.lmp:
Finally, let us use the minimize command to reduce the potential energy of the system:
These commands were covered in earlier tutorials and should already be familiar.
Then, let us densify the system to a target value of by imposing the shrinking of the simulation box at a constant rate. The dimension parallel to the CNT axis is maintained fixed because the CNT is periodic in that direction. Add the following commands to mixing.lmp:
The fix halt command is used to stop the box shrinkage once the target density is reached, and the other commands should be familiar from previous tutorials.
For the next stage of the simulation, we will use dump image to output images every 200 steps:
For the following , let us equilibrate the densified system in the constant-volume ensemble, and write the final state of the system in a file named mixing.data:
For visualization purposes, the atoms of the CNT group are moved to the center of the box using fix recenter. As the time progresses, the system density, , gradually converges toward the target value of g/cm3 (Fig. 48 a). Meanwhile, the total energy of the system initially evolves rapidly, reflecting the densification process, and then eventually stabilizes (Fig. 48 b). The final state is shown in Fig. 47.
3.8.2 Reaction templates
The REACTER protocol enables the modeling of chemical reactions using classical force fields. The user must provide a molecule template for the reactants, a molecule template for the products, and a reaction map file that provides an atom mapping between the two templates. The reaction map file also includes additional information, such as which atoms act as initiators for the reaction and which serve as edge atoms to connect the rest of a long polymer chain in the simulation.
There are three reactions to define: (1) the polymerization of two styrene monomers, (2) the addition of a styrene monomer to the end of a growing polymer chain, and (3) the linking of two polymer chains. Download the three files associated with each reaction. The first reaction uses the prefix ‘M-M’ for the pre-reaction template, post-reaction template, and reaction map file:
The second reaction uses the prefix ‘M-P’,
The third reaction uses the prefix ‘P-P’,
Here, the file names for each reaction use the abbreviation ‘M’ for monomer and ‘P’ for polymer.
Note 3.50.
The data stored in molecule templates include atom coordinates, partial charges, molecule IDs, atom types, and interaction types for bonds, angles, dihedrals and impropers. The map files contain information about the reaction. The first mandatory section of the map files begins with the keyword “InitiatorIDs” and lists the two atom IDs of the initiator atom pair in the pre-reacted molecule template. The second mandatory section begins with the keyword “Equivalences” and lists a one-to-one correspondence between atom IDs of the pre- and post-reacted templates. Some atoms in the pre-reacted template that are not reacting may have missing topology with respect to the simulation. For example, the pre-reacted template may contain an atom that, in the simulation, is currently connected to the rest of a long polymer chain. These are referred to as edge atoms, and are also specified in the map file in the “EdgeIDs” section.
3.8.3 Simulating the reaction
The first step, before simulating the reaction, is to import the previously generated configuration. Open the file named polymerize.lmp, which should contain the following lines:
Here, the read_data command is used to import the previously generated mixing.data file. All other commands have been introduced in earlier parts of the tutorial.
Then, let us import all six molecules templates using the molecule command:
In order to follow the evolution of the reaction with time, let us generate images of the system using dump image:
Let us use fix bond/react by adding the following line to polymerize.lmp:
With the stabilization keyword, the fix bond/react command will stabilize the atoms involved in the reaction using the fix nve/limit command with a maximum displacement of . The fix nve/limit command functions similar to fix nve, but restricts how far atoms can move in a single time step, even with very large forces. By default, each reaction is stabilized for 60 time steps. Each react keyword corresponds to a reaction, e.g., a transformation of mol1 into mol2. Implementation details about each reaction, such as the reaction distance cutoffs and the frequency with which to search for reaction sites, are also specified in this command.
Note 3.51.
The command fix bond/react creates several groups of atoms that are dynamically updated to track which atoms are being stabilized and which atoms are undergoing dynamics with the system-wide time integrator (here, fix nvt). When reaction stabilization is employed, there should not be a time integrator acting on the group all. Instead, the group of atoms not currently undergoing stabilization is named by appending ‘_REACT’ to the user-provided prefix.
Add the following commands to polymerize.lmp to carry out the dynamics using a Nosé-Hoover thermostat while ensuring that the CNT remains centered in the simulation box:
Here, the thermo_style custom command is used to print the cumulative reaction counts which are calculated by fix rxn and thus can be extracted from it. Run the simulation using LAMMPS. As the simulation progresses, polymer chains are observed forming (Fig. 49). During this reaction process, the temperature of the system remains well-controlled (Fig. 50 a), while the number of reactions, r, increases with time (Fig. 50 b).
Author Contributions
S.G. conceived and wrote the original online tutorials and underlying Sphinx documentation for lammpstutorials.github.io. C.A. tested the tutorials extensively, providing feedback that improved their clarity, accuracy, and usability. J.G. is the principal author of fix bond/react and type labels support in LAMMPS. He revised the tutorials to incorporate type labels and wrote Tutorial 8. A.K. developed the LAMMPS–GUI software and assisted in revising the tutorials for use with it. All authors participated in the revision and finalization of the manuscript.
Potentially Conflicting Interests
There are no conflicting interests to declare.
Funding Information
S.G. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement . A.K. acknowledges financial support by Sandia National Laboratories under POs 2149742 and 2407526. This work used Expanse at the San Diego Supercomputer Center through allocation MAT240081 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program.
Author Information
Appendix A Using LAMMPS–GUI
Note A.1.
For simplicity, these tutorials reference keyboard shortcuts based on the assignments for Linux and Windows. macOS users should use the “Command” key ( ) in place of the “Ctrl” key when using keyboard shortcuts.
A.1 Installation
Pre-compiled versions of LAMMPS–GUI are available for Linux, macOS, and Windows on the LAMMPS GitHub Release page [21]. The Linux version is provided in two formats: as compressed tar archive (.tar.gz) and as a Flatpak bundle [71]. The macOS version is distributed as a .dmg installer image, while the Windows version comes as an executable installer package.
A.1.1 Installing the Linux .tar.gz Package
Download the archive (e.g., LAMMPS-Linux-x86_64-GUI-22Jul2025.tar.gz) and unpack it. This will create a folder named LAMMPS–GUI containing the included commands, which can be launched directly using “./lammps-gui” or “./lmp”, for example. Adding this folder to the PATH environment variable will make these commands accessible from everywhere, without the need for the “./” prefix.
A.1.2 Installing the Linux Flatpak Bundle
You have to have Flatpak support installed on Linux machine to be able to use the Flatpak bundle. Download the bundle file (e.g., LAMMPS-Linux-x86_64-GUI-22Jul2025.flatpak) and then install it using the following command:
This will integrate LAMMPS–GUI into your desktop environment (e.g., GNOME, KDE, XFCE) where it should appear in the “Applications” menu under “Science”. Additionally, the “.lmp” file extension will be registered to launch LAMMPS–GUI when opening a file with this extension in the desktop’s file manager.
A.1.3 Installing the macOS Application Bundle
After downloading the macOS app bundle image file (e.g., LAMMPS-macOS-multiarch-GUI-22Jul2025.dmg), double-click on it. In the dialog that opens drag the LAMMPS–GUI app bundle into the Applications folder. To enable command-line access, follow the instructions in the README.txt file. These macOS app-bundles contain native executables for both, Intel and Apple CPUs.
After installation, you can launch LAMMPS–GUI from the Applications folder. Additionally, you can drag an input file onto the app or open files with the “.lmp” extension. Note that the LAMMPS–GUI app bundle is currently not cryptographically signed, so macOS may initially prevent it from launching. If this happens, you need to adjust the settings in the “Security & Privacy" system preferences dialog to allow access.
A.1.4 Installing the Windows package
Download the LAMMPS–GUI installer for Windows (e.g., LAMMPS-Win10-64bit-GUI-22Jul2025.exe). Windows may warn you that the file is from an unknown developer and was downloaded from the internet. This happens because neither the installer nor the LAMMPS–GUI application (or any other included applications) have been cryptographically signed. You will need to choose to keep the file, and when launching the installer, confirm that you want to run it despite the warning.
After installation, a new entry should appear in the Start menu. Additionally, the “.lmp” file extension should be registered with Windows File Explorer to open LAMMPS–GUI when opening a file with the “.lmp“ extension. The “lammps-gui” and “lmp” commands should also be available in the command-line.
A.2 Opening, Editing, and Saving Files
LAMMPS–GUI can be launched from the command-line, as explained above, where you can either launch it without arguments or provide one file name as an argument. All other arguments will be ignored. For example:
Files can also be opened from the “File” menu. You can select a file through a dialog and then open it. Additionally, a history of the last five opened files is maintained, with entries to open them directly. Finally, the Ctrl-O keyboard shortcut can also be used to open a file. When integrated into a desktop environment, it is also possible to open files with a “.lmp” extension or use drag-and-drop.
For the most part, the editor window behaves like other graphical editors. You can enter, delete, or copy and paste text. When entering text, a pop-up window will appear with possible completions after typing the first two characters of the first word in a line. You can navigate the highlighted options using the up and down arrow keys, and select a completion by pressing the Enter key or using the mouse. You can also continue typing, and the selection in the pop-up will be refined. For some commands, there will be completion pop-ups for their keywords or when a filename is expected, in which case, the pop-up will list files in the current folder.
As soon as LAMMPS–GUI recognizes a command, it applies syntax highlighting according to built-in categories. This can help detect typos, since those may cause LAMMPS–GUI not to recognize the syntax and thus not apply or partially apply the syntax highlighting. When you press the Tab key, the line will be reformatted. Consistent formatting can improve the readability of input files, especially long and complex ones.
If the file in the editor has unsaved changes, the word “*modified*” will appear in the window title. The current input buffer can be saved by selecting “Save” or “Save As…” from the “File” menu. You can also click the “Save” icon on the left side of the status bar, or use the Ctrl-S keyboard shortcut.
Note A.2.
When LAMMPS–GUI opens a file, it will switch the working directory to the folder that contains the input file. The same happens when saving to a different folder than the current working directory. The current working directory can be seen in the status bar at the bottom right. This is important to note because LAMMPS input files often require additional files for reading and may write output files (such as images, trajectory dumps, or averaged data files), which are typically expected to be in the same folder as the input file.
A.3 Running LAMMPS
From within the LAMMPS–GUI main window, LAMMPS can be started either from the «Run» menu by selecting the «Run LAMMPS from Editor Buffer» entry, using the keyboard shortcut Ctrl-Enter (Command-Enter on macOS), or by clicking the green «Run» button in the status bar. While LAMMPS is running, a message on the left side indicates that LAMMPS is active, along with the number of active threads. On the right side, a progress bar is displayed, showing the estimated progress of the current run or minimize command.
A.4 Creating Snapshot Images
Open the «Image Viewer» using either the «Create Image» option from the «Run» menu, the «Ctrl-I» keyboard shortcut, or click on the (right) palette button in the status bar. The image can be saved using the «Save As...» option from the «File» menu.
A.5 The Output Window
By default, when starting a run, the «Output» window opens to display the screen output of the running LAMMPS calculation. The text in the Output window is read-only and cannot be modified, but keyboard shortcuts for selecting and copying all or part of the text can be used to transfer it to another program: The keyboard shortcut «Ctrl-S» (or «Command-S» on macOS) can be used to save the Output buffer to a file. Additionally, the «Select All» and «Copy» functions, along with a «Save Log to File» option, are available through the context menu, which can be accessed by right-clicking within the text area of the «Output» window.
A.6 The Charts Window
By default, when starting a run, a «Charts» window opens to display a plot of the thermodynamic output from the LAMMPS calculation. From the «File» menu in the top-left corner, you can save an image of the currently displayed plot or export the data in various formats: plain text columns (for use with plotting tools like Gnuplot or XmGrace), CSV data (suitable for processing in Microsoft Excel, LibreOffice Calc, or Python with Pandas), or YAML (which can be imported into Python using PyYAML or Pandas). You can use the mouse to zoom in on the graph by holding the left button and dragging to select an area. To zoom out, right-click anywhere on the graph. You can reset the view by clicking the «lens» button located next to the data drop-down menu.
A.7 Preferences
The Preferences dialog allows customization of the behavior and appearance of LAMMPS–GUI. Among other options:
In the «General Settings» tab, the «Data update interval» setting allows you to define the time interval, in milliseconds, between data updates during a LAMMPS run. By default, the data for the «Charts» and «Output» windows is updated every 10 milliseconds. Set this to 100 milliseconds or more if LAMMPS–GUI consumes too many resources during a run. The «Charts update interval» controls the time interval between redrawing the plots in the «Charts» window, in milliseconds.
The «Accelerators» tab enables you to select an accelerator package for LAMMPS to use. Only settings supported by the LAMMPS library and local hardware are available. The «Number of threads» field allows you to set the maximum number of threads for accelerator packages that utilize threading.
The «Editor Settings» tab allows you to adjust the settings of the editor window. Select the «Auto-save on Run and Quit» option to automatically save changes made to the .lmp file upon closing LAMMPS–GUI.
See Ref. 16 for a full list of options.
Appendix B Running LAMMPS on the Command-Line without the GUI
LAMMPS can also be executed from the command-line on Linux, macOS, and Windows without using the GUI. This is the more common way to run LAMMPS. Both, the LAMMPS–GUI program and the LAMMPS command-line executable utilize the same LAMMPS library and thus no changes to the input file are required.
First, open a terminal or command-line prompt window and navigate to the directory containing the input.lmp file. Then execute:
where lmp is the command-line LAMMPS command.
For parallel execution with 4 processors (via OpenMP threads where supported by the OPENMP package), use:
Note B.1.
Running in parallel via MPI requires a specially compiled LAMMPS package and is not supported by the GUI. On supercomputers or HPC clusters, pre-compiled LAMMPS executables are typically provided by the facility’s user support team. For more information, please refer to the facility’s documentation or contact its user support staff.
See Ref. 72 for a complete description on how to run LAMMPS.